### v1

# Geodesic loops and orthogonal geodesic chords without self-intersections

Abstract.
We show that for a generic Riemannian metric on a compact manifold of dimension $n\geq 3$ all geodesic loops based at a fixed point have no self-intersections. We also show that for an open and dense subset of the space of Riemannian metrics on an $n$-disc with $n\geq 3$ and with a striclty convex boundary there are $n$ geometrically distinct orthogonal geodesic chords without self-intersections. We use a perturbation result for intersecting geodesic segments of the author [18] and a genericity statement due to Bettiol and Giambò [2] and existence results for orthogonal geodesic chords by Giambò, Giannoni, and Piccione [7].
Key words and phrases:
orthogonal geodesic chords without self-intersections, perturbation of Riemannian metrics, non-degenerate orthogonal geodesic chord, orthogonal-tangent geodesic chord, generic Riemannian metrics
2020 Mathematics Subject Classification:
53C22, 58E10
1. Results
The author has shown a perturbation result for geodesic segments in [18, Lem.3]. As a consequence we obtain for a point $p$ in a compact Riemannian manifold of dimension $n\geq 3$ with injectivity radius $\mathrm{inj}=4\eta>0$ the following statement: If $N\geq 2$ geometrically distinct geodesic segments $c_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,N$ parametrized by arc length intersect in the common midpoint $p=c_{j}(0)$ then one can perturb away the intersection of the geodesic segments. For a precise statement see Lemma 2.
[…]
Instead of considering closed resp. periodic geodesics one can study different boundary conditions for geodesic segments. In this paper we study first for points $p,q\in M$ geodesic segments $c:\mathcal{I}\longrightarrow M$ joining $p=c(0)$ and $q=c(1).$ Here we use as parameter space the unit interval $\mathcal{I}=[0,1]$ if $p\not=q$ and the circle $\mathcal{I}=S^{1}=\mathbb{R}/\mathbb{Z}=[0,1]/\{0,1\}$ in case of a geodesic loop, i.e. $p=q.$ We call the geodesic segment $c:\mathcal{I}\longrightarrow M$ simple, resp. without self-intersection, if the map $c:\mathcal{I}\longrightarrow M$ is injective. We call two geodesic segments $c_{1},c_{2}:\mathcal{I}\longrightarrow M$ geometrically distinct, if $c_{1}(\mathcal{I})\not=c_{2}(\mathcal{I}).$ We say that two geometrically distinct geodesic segments $c_{1},c_{2}:\mathcal{I}\longrightarrow M$ do not have an interior intersection, if $c_{1}((0,1))\cap c_{2}((0,1))=\emptyset.$ It follows from a general result by Bettiol and Giambò presented in  [2, Theorem 5.10] that for a generic Riemannian metric on a compact manifold $M$ the geodesic segments, resp. geodesic loops, are non-degenerate in the sense of Morse theory. We show in this paper that we can use the perturbation result Lemma 2 and the genericity result of [2, Theorem 5.10] to show the following result:
Theorem 1.
Let $M$ be a compact manifold of dimension $n\geq 3.$ Let $p,q\in M$ be points in $M,$ where we also allow $p=q.$ For a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all geodesic segments joining $p$ and $q$ are non-degenerate and simple, and geometrically distinct geodesic segments do not have an interior intersection.
Note that this also shows that for points $p,q$ for a generic metric there is no closed geodesic passing through $p$ and $q.$ It even shows that for a generic metric any geodesic through $p$ and $q$ does not go through $p$ resp. $q$ again. In the particular case $p=q$ it follows from Theorem 1 that all geodesic loops for a generic metric are prime, i.e. there is no shorter geodesic loop, which is a part of the given geodesic loop, cf. [16]. Hence for a generic metric for any two geodesics $c,d:[0,1]\longrightarrow M,c(0)=d(0)=p,c(1)=d(1)=q$ we have either $c((0,1))=d((0,1))$ or $c([0,1])\cap d([0,1])=\{p,q\}.$
As a consequence of Theorem 1 we obtain the following
Corollary 1.
Let $M$ be a simply-connected and compact manifold of dimension $n\geq 3$ and $p,q\in M.$ For a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all geodesic segments joining $p$ and $q$ are non-degenerate and simple, and geometrically distinct ones do not have interior intersection. For the number $N_{g}(p,q;t)$ of geodesic segments joining $p$ and $q$ of length $\leq t$ we obtain $\liminf_{t\to\infty}(N_{g}(p,q;t)/t)>0.$
This gives a quantitative version of Serre’s result [20, Ch.IV.,Sec.7, Prop.14] which also gives a geometrically satisfying result concerning geometrically distinct geodesic segments between two fixed points, cf. the discussion in the doctoral thesis by Mentges [16, Sec.1.2] or by Tanaka [21, Problem B]. Another boundary condition for geodesic segments is the following: Let $(M,g)$ be a compact Riemannian manifold with smooth boundary $\partial M$ and interior $\textrm{Int}(M)=M-\partial M$ then a geodesic chord $c:[0,1]\longrightarrow M$ is a geodesic segment with $c(0),c(1)\in\partial M$ and $c(t)\in\textrm{Int}(M)$ for all $t\in(0,1).$ It is called an orthogonal geodesic chord if $c$ meets the boundary $\partial M$ orthogonally. It is called an orthogonal-tangent geodesic chord if it is tangent to the boundary at $c(0)$ and orthogonal to the boundary at $c(1).$ As a consequence of the genericity result by Bettiol and Giambò [2] we show:
Theorem 2.
Let $M$ be a compact manifold of dimension $n\geq 3$ with smooth boundary $\partial M.$ For a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all orthogonal geodesic chords are non-degenerate and have no self-intersections. In addition they all do not intersect each other.
See Section 3 for the definition of non-degeneracy for orthogonal geodesic chords.
Lusternik and Schnirelmann [14] have shown that a bounded domain in $\mathbb{R}^{n},n\geq 2$ with a smooth and convex boundary has at least $n$ double normals, i.e. lines intersecting the boundary orthogonally. This was generalized for a Riemannian metric, Bos [4, p.431] showed that on an $n$-disc $D^{n}$ with a Riemannian metric whose boundary $\partial D^{n}$ is strictly convex there are $n$ geometrically distinct orthogonal geodesic chords. Using a generalization of this result due to Giambò, Giannoni, and Piccione [7, Theorem 1.1] we obtain from Theorem 2:
Theorem 3.
For $n\geq 3$ let $\mathcal{G}(D^{n})$ be the space of Riemannian metrics on the $n$-disc $D^{n}$ with the strong $C^{k}$-topology with $k\geq 2.$ The subset $\mathcal{G}_{1}=\mathcal{G}_{1}(D^{n})$ of Riemannian metrics on $D^{n}$ for which there is no orthogonal-tangent geodesic chord, is an open and non-empty subset of $\mathcal{G}(D^{n}).$ Then there is an open subset $\mathcal{G}_{2}\subset\mathcal{G}_{1}$ whose closure $\overline{\mathcal{G}}_{2}$ coincides with the closure $\overline{\mathcal{G}}_{1},$ i.e. $\overline{\mathcal{G}}_{1}=\overline{\mathcal{G}}_{2},$ such that for all $g\in\mathcal{G}_{2}$ there are $n$ geometrically distinct geodesic orthogonal chords which are non-degenerate and have no self-intersection and do not intersect each other.
Here we note that geometrically distinct geodesic chords cannot intersect at a boundary point. The boundary $\partial D^{n}$ is called strictly convex, if all principal curvatures $\kappa_{i},i=1,\ldots,n-1$ of the boundary with respect to the inward pointing unit normal, are positive, i.e. $\kappa_{i}>0,i=1,\ldots,n-1.$ In particular there is no geodesic in the disc touching the boundary $\partial D^{n}$ and hence there is no orthogonal-tangent geodesic chord. Therefore we obtain from Theorem 3 the following result in the spirit of Bos’ theorem [4, p.431]:
Corollary 2.
There is an open and dense subset $\mathcal{G}_{*}$ of the set $\mathcal{G}_{+}(D^{n})$ of Riemannian metrics on the $n$-dics $D^{n}$ of dimension $n\geq 3$ for which the boundary $\partial D^{n}$ is strictly convex such that for all Riemannian metrics $g\in\mathcal{G}_{*}$ there are at least $n$ geometrically distinct orthogonal geodesic chords without self-intersections. And these geodesic chords do not intersect each other.
Using a curve shortening construction called disc flow on surfaces Hass and Scott [11, Thm.4.2] have shown that a Riemannian $2$-disc with strongly convex boundary carries at least two geometrically distinct orthogonal geodesic chords without self-intersections. For the proof one can also use the curve shortening flow, as shown by Ko [13]. With a curve shortening flow Li [15] presents a proof of the above cited result by Lusternik and Schnirelmann about the existence of $n$ double normals of a strictly convex compact domain in $\mathbb{R}^{n}.$
2. Intersection of geodesic segments and perturbation
For the intersection points of geodesic segments we obtain the following result which is similar to the corresponding result for closed geodesics, cf. [18, Lem.2] resp. [2, Sec.2.3]. Recall from the Introduction that we denote a geodesic segment $c:\mathcal{I}\longrightarrow M$ joining $p=c(0),q=c(1)$ using the parameter space $\mathcal{I}=[0,1]$ if $p\not=q$ and $\mathcal{I}=S^{1}$ for $p=q.$
Lemma 1.
(a) Let $\gamma:\mathcal{I}\longrightarrow M$ be a geodesic segment, which is not the parametrization of a closed (periodic) geodesic of period $\leq 1.$ Then the set of interior double points, resp. interior self-intersection points
$DP_{1}(\gamma):=\{\gamma(t)\,;\,t\in(0,1),\#\gamma^{-1}(\gamma(t))\geq 2\}$
of the geodesic segment $\gamma$ is finite.
(b) Let $\gamma,\delta:[0,1]\longrightarrow M$ be geometrically distinct geodesic segments with $\gamma(0)=\delta(0)=p,\gamma(1)=\delta(1)=q.$ Then the set of interior intersection points
$I_{1}(\gamma,\delta):=\gamma((0,1))\cap\delta((0,1))$
is finite.
Remark 1.
(a) Note that in case (a) for a geodesic segment $\gamma:\mathcal{I}\longrightarrow M$ the initial point $c(0)$ is an interior self-intersection point, if there is a parameter $s\in(0,1)$ with $c(s)=c(0).$
(b) The geodesic segment $\gamma:\mathcal{I}\longrightarrow M$ has no interior self-intersection, resp. is called simple, if the map $\gamma$ is injective, i.e. if the set $DP_{1}(\gamma)$ of interior double points is empty. The geodesic segments $\gamma,\delta$ joining $p,q$ do not have interior intersection, if $I_{1}(\gamma,\delta)=\emptyset.$
Proof.
If $\mathrm{inj}>0$ is the injectivity radius of $(M,g),$ we conclude from $\gamma(t)=\gamma(s),t<s$ that $(s-t)L(\gamma)\geq\mathrm{inj}.$ This implies that for all $t\in[0,1]:\#\gamma^{-1}(\gamma(t))\leq L(\gamma)/\mathrm{inj}+1.$
(a) If the set $DP_{1}(\gamma)$ is not finite, then there exist sequences $s_{j}<t_{j}\in[0,1]$ converging to $s^{*},t^{*}\in[0,1]$ with $s_{j}\not=s^{*},t_{j}\not=t^{*},(t_{j}-s_{j})L(\gamma)\geq\mathrm{inj}$ for all $j,$ $(t^{*}-s^{*})L(\gamma)\geq\mathrm{inj},$ and $\gamma(s_{j})=\gamma(t_{j})$ for all $j\geq 1.$ For $p_{1}=\gamma(t^{*})=\gamma(s^{*})$ set $v=c^{\prime}(t^{*}),w=c^{\prime}(s^{*}).$ The geodesic segment $\gamma$ is parametrized proportional to arc length, hence $\|v\|=\|w\|.$ Since $c(t_{j})=\exp_{p}((t_{j}-t^{*})v)=c(s_{j})=\exp_{p}((s_{j}-s^{*})w)$ we obtain for sufficiently large $j$ with $|t_{j}-t^{*}|\|v\|,|s_{j}-s^{*}|\|w\|<\mathrm{inj}$ that $(t_{j}-t^{*})v=(s_{j}-s^{*})w$ holds, i.e. $v=\pm w.$ If $v=w$ then $\gamma(t)=\gamma(t+(t^{*}-s^{*}))$ for all $t,$ i.e. $\gamma$ is a closed geodesic, which is not prime. If $v=-w$ we conclude $\gamma((t^{*}-s^{*})/2-t)=\gamma((t^{*}-s^{*})/2+t)$ for all $t,$ showing that $\gamma$ is not smooth at $(t^{*}-s^{*})/2.$
(b) The argument is similar to (a): If the set $I_{1}(\gamma,\delta)$ is infinite then there are sequences $s_{j},t_{j}\in[0,1]$ converging to $s^{*},t^{*}$ with $s_{j}\not=s^{*},t_{j}\not=t^{*}$ and $\gamma(s_{j})=\delta(t_{j})$ for all $j\geq 1.$ Let $p=\gamma(s^{*})=\delta(t^{*})$ and $v=\gamma^{\prime}(s^{*}),w=\delta^{\prime}(t^{*}).$ Since $\gamma$ and $\delta$ are geometrically distinct, we have $v/\|v\|\not=\pm w/\|w\|.$ Since $\gamma(t_{j})=\exp_{p}((t_{j}-t^{*})v)=\delta(s_{j})=\exp_{p}((s_{j}-s^{*})w)$ we conclude for sufficiently large $j$ with $|t_{j}-t^{*}|\|v\|,|s_{j}-s^{*}|\|w\|<\mathrm{inj}$ that $(t_{j}-t^{*})v=(s_{j}-s^{*})w$ holds, i.e. $v/\|v\|=\pm w/\|w\|,$ which is a contradiction. ∎
We obtain from the perturbation result [18, Lem.3]:
Lemma 2.
Let $(M,g)$ be a compact Riemannian manifold of dimension $n\geq 3$ and $\eta>0$ be a positive number such that the injectivity radius $\mathrm{inj}$ satisfies $\mathrm{inj}\geq 4\eta.$ For a point $p\in M$ let $c_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,k$ be $k\geq 2$ geometrically distinct geodesic segments parametrized by arc length with common midpoint $p=c_{j}(0),j=1,\ldots,k.$
Then there is in any neighborhood of the Riemannian metric $g$ with respect to the strong $C^{r}$-topology with $r\geq 2$ a Riemannian metric $\overline{g}$ coinciding with $g$ outside the geodesic ball $B_{p}(\eta)$ of radius $\eta$ centered at $p$ carrying geodesic segments $\overline{c}_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,k$ parametrized by arc length which coincide with $c_{j}$ outside $B_{p}(\eta).$ These geodesic segments are simple and do not intersect each other.
A similar proof as in Lemma [18, Lem.3] shows the following
Lemma 3.
Let $c:[0,1]\longrightarrow M$ be a geodesic loop with $p=c(0)=c(1)$ of the Riemannian metric $g$ on the differentiable manifold $M$ with $v=c^{\prime}(0)\in T_{p}M,w=c^{\prime}(1)\in T_{p}M.$ Then there exists a continuous one parameter perturbation $g_{s},s\in[0,s_{1}]\in\mathcal{G}(M)$ in the space of Riemannian metrics with the strong $C^{k}$-topology with $k\geq 2$ of the metric $g=g_{0}$ and continuous functions $s\in[0,s_{1}]\mapsto v(s),w(s)\in T_{p}M$ such that $c_{s}:[0,1]\longrightarrow M$ is a geodesic loop with $p=c(0)$ of the Riemannian metric $g_{s}$ with $v(s)=c_{s}^{\prime}(0),w(s)=c_{s}^{\prime}(1)$ and $v(s),v$ resp. $w(s),w$ are linearly independent for $s>0.$ The metric $g_{s}$ agrees with $g$ outside an arbitrarily small geodesic ball around $p.$
3. Geodesics with product boundary conditions
The following concept is introduced in [2, Sec.4], cf. also [10]. Let $M$ be a differentiable manifold. Let $\mathcal{P}\subset M\times M$ be a compact submanifold. This is called a general boundary condition on $M$ if for some Riemannian metric $g$ on $M$ the pseudo-Riemannian metric $\overline{g}=g\oplus(-g)$ on $M\times M$ induces a non-degenerate pseudo-Riemannian metric on $\mathcal{P},$ cf.  [2, Def.4.2]. It is shown in [2, Prop.4.1] that the subset $\mathcal{A}_{\mathcal{P}}\subset\mathcal{G}(M)$ of the set of Riemannian metrics on $M$ for which the restriction of the pseudo Riemannian metric $\overline{g}$ onto the submanifold $\mathcal{P}$ is non-degenerate, is an open subset. The second fundamental form of $\mathcal{P}$ in the normal direction $\eta\in T\mathcal{P}^{\perp}$ is given by
$S^{\mathcal{P}}_{\eta}(v,w)=\overline{g}\left(\nabla_{v}^{\overline{g}}\overline{w},\eta\right),$
cf. [2, p.346]. Here $\overline{\nabla}$ is the Levi-Civita connection on $M\times M$ with metric $\overline{g}.$ The vector field $\overline{w}$ is a smooth extension of $w$ to a smooth vector field along $\mathcal{P}.$ Since the restriction of $\overline{g}$ to $\mathcal{P}$ is non-degenerate we can identify the symmetric bilinear form with the self-adjoint shape operator $(S_{\eta}^{\mathcal{P}})_{p}:T_{p}\mathcal{P}\longrightarrow T_{p}\mathcal{P}$ with $\overline{g}((S_{\eta}^{\mathcal{P}})_{p}v,w)=S_{\eta}^{\mathcal{P}}(v,w)$ for all $v,w\in T_{p}\mathcal{P}.$ Let $H^{1}(I,M)$ be the set of all curves of Sobolev class $H^{1}$ in $M,$ i.e. of absolutely continuous curves with finite energy. Then the subspace
$\Omega_{\mathcal{P}}(M)=\left\{\gamma\in H^{1}(I;M)\,;\,(\gamma(0),\gamma(1))\in\mathcal{P}\right\}$
is a Hilbert manifold. The tangent space $T_{\gamma}\Omega_{\mathcal{P}}(M)$ can be identified with the completion of the space of smooth vector fields $v$ along $\gamma$ with $(\gamma(0),\gamma(1))\in T_{(\gamma(0),\gamma(1))}\mathcal{P}$ with respect to the induced metric $g_{1}=\langle.,.\rangle_{1}.$ It is defined on $T_{\gamma}\Omega_{\mathcal{P}}(M)$ by
[…]
A geodesic $\gamma:I\longrightarrow M$ is called a $(g,\mathcal{P})-$geodesic if $(\gamma^{\prime}(0),\gamma^{\prime}(1))\in T_{(\gamma(0),\gamma(1))}\mathcal{P}^{\perp}.$ Here the orthogonality is with respect to the metric $\overline{g}.$ It is shown in [2, Sec. 5] that a curve $\gamma$ is a critical point of the energy functional $E_{g}:\Omega_{\mathcal{P}}(M)\longrightarrow\mathbb{R},$
[…]
if and only if it is a $(g,\mathcal{P})$-geodesic. The kernel of the index form is the space of Jacobi fields $Y\in T_{\gamma}\Omega_{\mathcal{P}}(M)$ such that
(2) $\left(\nabla Y(0),\nabla Y(1)\right)+S_{(\gamma^{\prime}(0),\gamma^{\prime}(1))}(Y(0),Y(1))\in T_{(\gamma(0),\gamma(1))}\mathcal{P}^{\perp}.$
[…]
A general boundary condition is called admissible if for every $g_{0}\in\mathcal{A}_{\mathcal{P}}$ there exists an open neighborhood $\mathcal{V}$ of $g_{0}$ in $\mathcal{A}_{\mathcal{P}}$ and a positive real number $a>0$ such that for all $g\in\mathcal{V}$ and all non-trivial $(g,\mathcal{P})$-geodesics $\gamma$ we have: $L_{g}(\gamma)\geq a,$ cf. [2, Def.4.5]. Here a non-trivial $(g,\mathcal{P})$ geodesic is a non-constant curve.
We restrict ourselves to the following case:
Definition 1.
Let $(M,g)$ be a complete Riemannian manifold of dimension $n$ and $N_{0},N_{1}\subset M$ connected and compact submanifolds (without boundary) of dimension $\dim N_{i}=n_{i}\in\{0,1,\ldots,n-1\},i=0,1.$ Then the product boundary condition is given as general boundary condition with $\mathcal{P}=N_{0}\times N_{1}.$
In particular a $(g,N_{0}\times N_{1})$-geodesic is a geodesic $c:[0,1]\longrightarrow M$ with $c(l)\in N_{l},c^{\prime}(l)\in T_{c(l)}^{\perp}N_{l},l=0,1.$ The product metric $g\oplus(-g)$ on $N_{0}\times N_{1}$ is non-degenerate, hence a product boundary condition is always non-degenerate, as defined in [2, Sec.4]. In case $\mathcal{P}=N\times N$ the $(g,N\times N)$-Jacobi fields are Jacobi fields $Y\in T_{\gamma}\Omega_{N\times N}$ with
(3) $\nabla Y(l)+S_{\gamma(l)}(Y(l))\in T_{\gamma(l)}N^{\perp}\,,l=0,1\,,$
they form the kernel of the index form. The $(g,N\times N)$-geodesic $\gamma$ is non-degenerate if the kernel is trivial, i.e. if there is no non-trivial $(g,N\times N)$-Jacobi field. Here $S_{x}$ denotes the shape operator of the submanifold $N\subset M$ with respect to a unit normal field $\nu,$ i.e.
$S_{x}(v)=\nabla_{v}\nu\,.$
And $\nabla$ is the Levi-Civita connection of $(M,g).$ If $N_{0}=\{p\},N_{1}=\{q\},$ then a $(g,\{p\}\times\{q\})$-Jacobi field along a $(g,\{p\}\times\{q\})$-geodesic segment $\gamma:[0,1]\longrightarrow M$ is a Jacobi field $Y$ vanishing at the end points, i.e. $Y(0)=0,Y(1)=0.$ Hence $q$ is not a conjugate point for $p$ along $\gamma$ if and only if there is no $(g,\{p\}\times\{q\})$-Jacobi field.
Lemma 4.
(a) Let $(M,g)$ be a compact Riemannian manifold (without boundary) of dimension $n\geq 2.$ If $p,q\in M$ are two points, where we also allow $p=q,$ then $\{p\}\times\{q\}$ is an admissible product boundary condition.
(b) If $(M,g)$ is a compact Riemannian manifold of dimension $n\geq 2$ with smooth boundary $\partial M$ then $\partial M\times\partial M$ is an admissible product boundary condition.
(c) If $(M,g)$ is a compact Riemannian manifold (without boundary) of dimension $n\geq 2$ and if $N\subset M$ is a compact submanifold of codimension one then $N\times N$ is an admissible product boundary condition.
Remark 2.
In case (b) we can consider the compact manifold $M$ with smooth boundary $\partial M$ endowed with a Riemannian metric $g$ as subset of a complete Riemannian manifold $(M_{1},g_{1}).$
Proof.
(a) We can choose $a=d(p,q)/2$ if $p\not=q,$ and $a=\mathrm{inj}/2$ for $p=q,$ here $\mathrm{inj}$ is the injectivity radius. Here we use that the distance as well as the injectivity radius are continuous functions on $\mathcal{G}^{r}(M)$ for $r\geq 2.$
(b) and (c). Here the normal exponential map of the submanifold $\partial M$ resp. $N$ is defined $\exp^{\perp}:\nu(\partial M)\longrightarrow M,$ resp. $\exp^{\perp}:\nu(N)\longrightarrow M.$ Here $\nu(\partial M)$ is the subset of the normal bundle of $\partial M$ of normal vectors pointing inward and $\nu(N)$ is the normal bundle of $N\subset M.$ Then there is a positive and continuous function $g\in\mathcal{G}^{r}(M)\mapsto\mathrm{inj}^{\perp}(g)$ such that the restriction onto the normal vectors of length $\leq\mathrm{inj}^{\perp}$ is a diffeomorphism onto its image. Any non-trivial $(g,\partial M\times\partial M)$-geodesic resp. $(g,N\times N)$-geodesic $\gamma:[0,1]\longrightarrow M$ satisfies $L_{g}(\gamma)\geq 2\mathrm{inj}^{\perp}(g).$
We will discuss the following examples in detail: If for some points $p,q\in M$ we have $N_{0}=\{p\},N_{1}=\{q\},$ then a $(g,\{p\}\times\{q\})$-geodesic is a geodesic segment joining $p$ and $q.$ Here we also allow the case $p=q,$ i.e. the case of geodesic loops based at $p.$ Another important case is $N_{0}=N_{1}=\partial N$ is a compact smooth boundary of an open subset $N$, then orthogonal geodesic chords are $(g,\partial{N}\times\partial{N})$-geodesics. And for a compact hypersurface $N\subset M$ i.e. a compact hypersurface $N$ we consider $(g,N\times N)$-geodesics, i.e. geodesics $c:[0,1]\longrightarrow M$ with $c(l)\in N,c^{\prime}(l)\in T_{c(l)}^{\perp}N,l=0,1.$
Lemma 5.
Let $M$ be a compact manifold of dimension $n\geq 2$ with smooth boundary $\partial M.$ Then there exists a compact manifold $M_{1}$ without boundary, such that $M$ is an open subset of $M_{1}$ and $\partial M$ is a smooth submanifold of $M_{1}.$
Proof.
It is a standard construction that one can add a cylindrical neck to the manifold $M$ with smooth boundary. I.e. there exists a Riemannian manifold $M_{0}$ with metric $g_{0}$ such that $M$ is an open submanifold and $g$ is the restriction of $g_{0}$ to $M.$ And for some positive $\epsilon>0$ the normal exponential map of the hypersurface $\partial M$ defines an injective immersion $\phi:(t,x)\in[-\epsilon,\epsilon]\times\partial M\longmapsto M_{0}.$ The Riemannian metric then has the form $dt^{2}+h_{t}$ for a smooth family $t\in[-\epsilon,\epsilon]\mapsto h_{t}\in\mathcal{G}(\partial M)$ of Riemannian metrics $h_{t}$ on $\partial M$ and $h_{t}=h_{1}$ for all $t\in[2/3,1].$ Hence the set $\phi([2/3,1]\times\partial M)$ is the cylindrical end, which is isometric to the Riemannian product $[2/3,1]\times(\partial M,h_{1}).$ This construction can be found for example in [8, Section 5].
Then one can take two copies of $(M_{0},g_{0})$ identified along the boundary $\partial M_{0},$ hence we obtain a compact Riemannian manifold $(M_{1},g_{1})$ without boundary such that $\partial M_{0}$ is a totally geodesic submanifold. ∎
Remark 3.
In the setting of Lemma 5 let $\phi:\mathcal{G}(M_{1})\longrightarrow\mathcal{G}(M)$ be the map defined by restricting the Riemannian metric $g$ on $M_{1}$ to the submanifold $M,$ i.e. $\phi(g)=g|M.$ The following statements are an immediate consequence of the definition of the strong topology, in particular of the definition of a strong basic neighborhood cf. [12, Section 2.1] applied to the space of Riemannian metrics on $M$ resp. $M_{1}.$ After fixing a Riemannian metric the space $\mathcal{G}(M)$ of Riemannian metrics of class $C^{r},r\geq 2$ on the manifold $M$ endowed with with the strong topology can be seen as an open subset in the Banach space of $C^{r}$-sections of the bundle of symmetric bilinear forms on the manifold $M.$ If $\mathcal{U}$ is an open and dense subset of $\mathcal{G}(M_{1})$ then $\phi(\mathcal{U})$ is an open and dense subset of $\mathcal{G}(M).$ And if $\mathcal{V}$ is a residual subset of $\mathcal{G}(M_{1}),$ then the restriction $\phi(\mathcal{V})$ is a residual subset of $\mathcal{G}(M).$
Lemma 6.
Let $L>0.$
(a) Let $(M,g)$ be a compact Riemannian manifold of dimension $n\geq 3$ and $p,q\in M.$ Then there is an open and dense set $\mathcal{G}^{*}[L]$ of Riemannian metrics such that for any $g\in\mathcal{G}^{*}[L]$ no geodesic segment of length $\leq L$ joining $p$ and $q$ is part of a closed geodesic.
(b) Let $(M,g)$ a compact Riemannian manifold of dimension $n\geq 2$ and let $N\subset M$ be a compact submanifold of codimension $1$ Then there is an open and dense set $\mathcal{G}^{*}[L]$ of Riemannian metrics on $M$ such that for any $g\in\mathcal{G}^{*}[L]$ no $(g,N\times N)$-geodesic of length $\leq L$ is part of a closed geodesic.
Proof.
Assume that $\gamma:S^{1}=[0,1]/\{0,1\}\longrightarrow M$ is a closed geodesic. We show that the set $A(\gamma):=\{t\in S^{1};\gamma^{\prime}(t)\in T_{c(t)}^{\perp}N_{l}\},l=0,1$ is finite. Since the boundary condition is admissible there is a positive $a>0$ such that for $s,t\in A,s\not=t$ we have $|t-s|L(\gamma)\geq a.$ This implies that $A$ is a finite set.
For $L>0$ let $\mathcal{G}^{*}[L]$ be the set of Riemannian metrics for which all closed geodesics with energy $\leq L$ are non-degenerate and simple, cf. [18, Thm.1] such that no restriction onto a subinterval defines - up to linear reparametrization - a geodesic segment between $p$ and $q$ in case (a) resp. a $(g,N\times N)$-geodesic in case (b). This set is open, we have to show that it is a dense subset of $\mathcal{G}(M).$ A metric $g\in\mathcal{G}^{*}[L]$ has only finitely many geometrically distinct prime closed geodesics $c_{j}:S^{1}\longrightarrow M,j=1\ldots,k$ of energy $\leq L$ and there is an open subset $\mathcal{U}\subset\mathcal{G}^{*}[L]$ and continuous functions $g\in\mathcal{U}\mapsto v_{j}(g)\in TM$ such that the curve $t\in\mathbb{R}\mapsto c_{j}(t)=\Phi_{g}^{t}(v_{j}(g)),t\in S^{1}$ is a closed geodesic of $g$ and there are - up to geometric equivalence - no further closed geodesics of length $\leq L$ of the metric $g,$ cf. [1, p.12-13], resp. the Proof of Part (b) of Theorem 4. Here $\Phi^{t}_{g}:TM\longrightarrow TM,t\in\mathbb{R}$ denotes the geodesic flow, i.e. for $v\in T_{p}M,t\in\mathbb{R}\mapsto\Phi_{g}^{t}(v)\in TM$ is the periodic orbit in the tangent bundle whose projection on the manifold is the geodesic $c:\mathbb{R}\longrightarrow M$ with initial point $p=c(0)$ and initial direction $v=c^{\prime}(0).$
Hence the set $\widetilde{A}=\bigcup_{i=1}^{k}\{c_{j}(t)\,;\,t\in A(c_{j})\}$ is finite.
(a) We can perturb the metric in a sufficiently small neighborhood of the points $p$ and $q$ using Lemma 2 such that the perturbed closed geodesics of length $\leq L$ do not meet $p$ and $q.$
(b) We perturb the metric in a sufficiently small neighborhood of $p=c_{j}(t)\in\widetilde{A},t\in A(c_{j})$ such that $c_{j}$ remains a closed geodesic but $c_{j}^{\prime}(t)\not\in T_{c_{j}(t)}^{\perp}N_{l}$ with respect to the perturbed metric.
This is possible as the following Lemma shows. ∎
Lemma 7.
Let $D^{n-1}_{2\eta}=\{x\in\mathbb{R}^{n-1}\,;\,\|x\|\leq 2\eta\}$ be the Euclidean $(n-1)$-disc of radius $2\eta$ for $\eta>0.$ On $(t,x)=(t,x_{1},\ldots,x_{n-1})\in[-2\eta,2\eta]\times D^{n-1}_{2\eta}$ a Riemannian metric is given by
$dt^{2}+\sum_{i,j=1}^{n-1}g_{ij}(t,x)dx_{i}dx_{j},$
hence $\gamma(t)=(t,0,\ldots,0),t\in[-2\eta,2\eta]$ is a geodesic of $g$ parametrized by arc length orthogonal at $p=\gamma(0)=(0,\ldots,0)$ to the hypersurface $\Sigma:=\{0\}\times D^{n-1}_{2\eta}.$ Then in any neighborhood of the metric $g$ there is a metric $\overline{g}$ for which $\gamma$ is still a geodesic parametrized by arc length and which is not orthogonal to $\Sigma$ at $p$ with respect to $\overline{g}.$
Proof.
Choose a smooth bump function $f:[-2\eta,2\eta]\longrightarrow[0,1]$ with $f(t)=0$ for $|t|\geq\eta,$ $f(t)=1$ for $|t|\leq\eta/2.$ Let $s=t+\epsilon f(t)x_{1}.$ For sufficiently small $\epsilon>0$ with inverse function $t_{\epsilon}(s,x_{1})$ in the coordinates $s,x_{1},\ldots,x_{n-1}$ the metric is of the form
$\frac{1}{(1-\epsilon f^{\prime}(t))^{2}}\left(ds^{2}+2\epsilon f(t)dsdx_{1}+\epsilon^{2}f^{2}(t)dx_{1}^{2}\right)+\sum_{i,j=1}^{n-1}g_{ij}(t_{\epsilon}(s,x_{1}),x_{1},\ldots,x_{n-1})dx_{i}dx_{j}\,.$
Then for sufficiently small $\epsilon>0$ the metric $\overline{g}$ given by
$ds^{2}+\sum_{i,j=1}^{n-1}g_{ij}(t_{\epsilon}(s,x_{1}),x_{1},\ldots,x_{n-1})dx_{i}dx_{j}$
is sufficently close go $g$ in the $C^{r}$-topology for $r\geq 2$ and the the curve $x_{1}=\ldots=x_{n-1}=0$ coincides with $\gamma$ and is a geodesic of $\overline{g}$ parametrized by arc length. At the point $p=\gamma(0)$ the curve is orthogonal to the hypersurface $s=0$ but not to the hypersurface $\Sigma=\{t=0\}.$
Theorem 4.
Let $(M,g)$ be a compact Riemannian manifold of dimension $n\geq 3.$
(a) Let $M$ be a compact differentiable manifold of dimension $n\geq 3.$ Let $p,q\in M.$ Then for a $C^{k}$-generic Riemannian metric with $k\geq 2$ on $M$ all geodesic segments between $p$ and $q$ are non-degenerate and simple, and any geometrically distinct geodesic segments between $p$ and $q$ do not have interior intersection.
(b) Let $M$ be a compact differentiable manifold of dimension $n\geq 3$ with a smooth boundary $\partial M.$ Then for a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all orthogonal geodesic chords are non-degenerate and simple, and any two geometrically distinct orthogonal geodesic chords do not intersect.
(c) Let $M$ be a compact differentiable manifold (without boundary) of dimension $n\geq 3$ and let $N\subset M$ be a compact submanifold of codimension $1$ Then for a $C^{k}$-generic metric on $M$ with $k\geq 2$ all $(g,N\times N)$-geodesics are non-degenerate and simple, and any two geometrically distinct $(g,N\times N)$-geodesics do not intersect.
Proof.
We first consider the three cases (a), (b), (c) at the same time, i.e. we set $\mathcal{P}=N_{0}\times N_{1}=\{p\}\times\{q\},$ or $\mathcal{P}=\partial{M}\times\partial{M},$ or $\mathcal{P}=N\times N.$
It follows from Lemma 6 and from the result [2, Thm.5.1], cf. also [2, p.347], that for $L>0$ the set $\mathcal{G}_{\mathcal{P}}(M)[L]$ of metrics $g\in\mathcal{G}_{\mathcal{P}}(M)$ for which all $(g,\mathcal{P})$-geodesics of length $\leq L$ are nondegenerate and are not part of a closed geodesic is an open and dense subset of $\mathcal{G}_{\mathcal{P}}=\mathcal{G}_{\mathcal{P}}(M).$ This holds since in all three cases (a),(b),(c) the boundary conditions $\mathcal{P}$ are non-degenerate and admissible, cf. Lemma 4.
Here one has to note that in case (b) one has to consider the differentiable manifold $M$ to be a subset of a compact manifold $M_{1}$ without boundary, cf. Lemma 5. Then one can apply [2, Thm.5.1] and conclude that the set $\mathcal{G}_{\partial M\times\partial M}(M_{1})[L]$ of metrics $g\in\mathcal{G}(M_{1})$ for which all $(g,\partial M\times\partial M)$-geodesics of length $\leq L$ are non-degenerate is an open and dense subset of $\mathcal{G}_{\partial M\times\partial M}(M_{1}).$ Then the set $\mathcal{G}_{\partial M\times\partial M}(M)$ contains all restrictions $g|M$ of metrics $g\in\mathcal{G}_{\partial M\times\partial M}(M_{2})[L],$ and is therefore also dense in $\mathcal{G}(M),$ cf. Remark 3.
For any $L>0$ we define the set $\mathcal{G}_{\mathcal{P}}^{*}(M)[L]$ of metrics $g\in\mathcal{G}^{L}(M)$ for which all $(g,\mathcal{P})$-geodesics of length $\leq L$ are nondegenerate and simple and such that for any pair $\gamma,\delta$ of geometrically distinct $(g,\mathcal{P})$-geodesics of length $\leq L$ in case of $\mathcal{P}=\{p\}\times\{q\}$ the interior intersection set $I_{1}(\gamma,\delta)$ is empty, and in the other two cases the intersection $\gamma(I)\cap\delta(I)=\emptyset.$
We conclude from [10, Thm.2.4] that the energy functional $E_{g}:c\in\Omega_{\mathcal{P}}(M)\longmapsto E_{g}(c)\in\mathbb{R}$ satisfies the Palais-Smale condition. We can apply this result since $M$ is compact, hence any Riemannian metric on $M$ is complete.
Hence for $g\in\mathcal{G}_{\mathcal{P}}(M)[L]$ the restriction $E_{g}:\Omega_{\mathcal{P}}(M)\cap\{E\leq L\}\longrightarrow\mathbb{R}$ is a Morse function with only finitely many critical points, i.e. there are only finitely many geometrically distinct $(g,\mathcal{P})$-geodesics $c_{j},j=1,\ldots,k$ of energy $\leq L.$
The self-adjoint operator $H_{c}$ representing the hessian $d^{2}E_{g}(c)$ at a $(g,\mathcal{P})$-geodesic $c$ is of the form $H_{c}=\mathbbm{1}+k_{c},$ i.e. the sum of the identity and a compact operator $k_{c}.$ Since 00 is not an eigenvalue the inverse $H_{c}^{-1}$ is a bounded linear operator. Therefore we can use the implicit function theorem for Banach manifolds, cf. [5, Thm.15.1] and conclude: There is a neigbhourhood $\mathcal{U}$ of $g$ and continuous functions $g\in\mathcal{U}\longmapsto c_{j}[g]\in\Omega_{\mathcal{P}}(M),j=1,\ldots,k$ such that $c_{j}[g]$ is a $(g,\mathcal{P})$-geodesic of $g\in\mathcal{U}.$
One can choose the open neighborhhod $\mathcal{U}$ sufficiently small such that there are no further $(g,\mathcal{P})$-geodesics of $g$ with $E_{g}(\gamma)\leq L,$ cf. the argument in [1, p.12-13].
Since the geodesics $c_{j}$ are not part of a closed geodesic we conclude from Lemma 1: The union $DIP_{1}(L)$ of the double points $DP_{1}(c_{j}),j=1,\ldots,k$ and the intersection points $I(c_{j},c_{l}),j\not=l,j,k\in\{1,\ldots,k\}$ is finite.
Let $x\in DIP_{1}(L).$ In cases (b)(c) it follows that if $x=c_{j}(t)$ for some $j=1\ldots,k$ then $t\in(0,1).$ In case (b) a geodesic chord $c_{j}:[0,1]\longrightarrow M$ satisfies $c_{j}(t)\not\in\partial M$ for all $t\in(0,1),$ hence $DIP_{1}(L)\cap\partial M=\emptyset.$
In case (c) if $x\in DIP_{1}(L)$ satisfies $x=c_{j}(0)=c_{j}(t)$ for some $t\in(0,1]$ then $c_{j}$ lies on a closed geodesic which is not possible by assumption. If $x=c_{j}(0)=c_{k}(t)$ for $j\not=k$ then $t\not\in\{0,1\}$ since otherwise the geodesics $c_{j},c_{k}$ lie on the same geodesic line $\gamma:\mathbb{R}\longrightarrow M$ since $N$ has codimension one. If $x=c_{j}(0)=c_{i}(t)$ for $i,j=1,\ldots,k,0<t<1$ then we can use the perturbation result Lemma 2 and obtain a Riemannian metric for which the geodesic segment $c_{j}$ in the neighborhood remains unchanged and the geodesic segment $c_{k}$ in the neighboorhood is perturbed such that $c_{j}$ and $c_{k}$ do not intersect in the neighborhood of $x.$
Now we consider case (a).
We first assume $p=q.$ For $L>0$ for a Riemannian metric $g\in\mathcal{G}_{\{p\}\times\{p\}}(M)[L]$ there are only finitely many geometrically distinct and non-degenerate geodesic loops $\gamma_{1},\ldots,\gamma_{N}$ of length $\leq L.$ Assume that - after a possible renumbering - the geodesic loops $\gamma_{1},\ldots,\gamma_{m},m\leq N$ are the prime geodesic loops of length $\leq L.$ Here a geodesic loop $\gamma:[0,1]/\{0,1\}\longrightarrow M$ is prime, if there is no $t_{1}\in(0,1)$ with $p=c(t_{1}).$ Let $v_{j,l}=\gamma_{j}^{\prime}(l),j=1,\ldots,m,l\in\{0,1\}$ be the sequence of tangent vectors of the prime geodesic loops at $p.$ We use the perturbation result Lemma 3 which shows that there is an open and dense subset $\mathcal{G}_{p}[L]\subset\mathcal{G}_{\{p\}\times\{p\}}(M)[L]$ for which all geodesic loops at $p$ of length $\leq L$ are prime. Note here that if $m<N,$ then there the geodesic loops $\gamma_{m+1},\ldots,\gamma_{N}$ are not prime, but their prime parts are loops $\gamma_{j},1\leq j\leq m.$ Since after the perturbation these prime geodesic loops cannot be concatenated to produce a non-prime geodesic loop the perturbed loops $\gamma_{j}[g],j\geq m+1$ have to be prime, too.
In case $p\not=q$ we consider the space $\mathcal{G}_{p,q}[L]=\mathcal{G}_{p}[L]\cap\mathcal{G}_{q}[L]\cap\mathcal{G}_{\{p\}\times\{q\}}(M)[L].$ Hence for $g\in\mathcal{G}_{p,q}[L]$ all geodesic loops at $p$ and $q$ of length $\leq L$ are non-degenerate and prime and there are only finitely many geometrically distinct geodesic loops at $p$ and $q.$ And all geodesic segments joining $p$ and $q$ of length $\leq L$ are non-degenerate and there are only finitely many geometrically distinct ones. Hence the set of tangent vectors of geodesic loops at $p$ resp. $q$ of length $\leq L$ as well as the set of tangent vectors of prime geodesic segments joining $p$ and $q$ of length $\leq L$ at $p$ and $q$ is finite. Therefore we can use the perturbation lemma 3 a finite number of times to obtain a metric for which all geometric segments joining $p$ and $q$ are prime. Here the observation is that a geodesic segment between $p$ and $q$ which is not prime contains either distinct prime geodesic segments joining $p$ and $q$ or contains prime geodesic segments and geodesic loops at $p$ or at $q.$
Therefore this argument shows that for points $p,q$ – where here we can also allow $p=q$ – the set $\mathcal{G}^{\prime}_{\{p\}\times\{q\}}[L]$ of Riemannian metrics for which all geodesic segments of length $\leq L$ joining $p$ and $q$ are non-degenerate and prime is an open and dense subset. If now $x\in DIP_{1}(L),x\not\in\{p,q\}$ we can choose $\eta\in(0,\mathrm{inj}/3)$ such that the following holds: If a geodesic $c_{j}$ for some $j=1,\ldots,k$ enters the geodesic ball $B_{p}(2\eta)$ around $p\in DIP(L)$ with radius $2\eta,$ i.e. $c_{j}(s)\in B_{p}(2\eta),$ then there is $s_{1}$ with $c_{j}(s_{1})=p$ and $d(c_{j}(s),c_{j}(s_{1}))=|s-s_{1}|\|c_{j}^{\prime}\|\leq 2\eta.$ Hence one can use the perturbation result [18, Lem.3] resp. Lemma 2 and perturb the metric in the ball $D_{p}(2\eta)$ for all finitely many points in $DIP_{1}(L)$ such that we obtain a metric $g_{1}\in\mathcal{G}_{\{p\}\times\{q\}}^{*}(M)[L]$ in any arbitrary small neighborhood of the original metric. ∎
4. Proofs of the statements in section 1
Theorem 1 is Part (a) of Theorem 4.
Proof of Corollary 1.
Since $M$ is simply-connected and compact there is a number $l\in\{2,\ldots,n-1\},$ such that the sequence $b_{2lk}=\dim H_{2lk}\left(\Omega_{\{p\}\times\{q\}}(M);\mathbb{Q}\right),k\geq 1$ satisfies $b_{2lk}\geq 1.$ It was shown by Gromov th,at there is a constant $C(g)>0$ such that the critical value $\mathrm{cr}(w_{k})$ of a non-trivial homology class $w_{k}\in H_{2lk}\left(\Omega_{\{p\}\times\{q\}}(M);\mathbb{Q}\right)$ satisfies $\mathrm{cr}(w_{k})\leq C(g)k$ for all $k\geq 1,$ cf. Gromov [9], [17, Thm.5.1]. By Theorem 1 we obtain for a generic metric a sequence $c_{k}$ of simple and non-degnerate geodesic segments joining $p$ and $q$ of index $2lk$ and length $\leq C(g)k.$ This follows from the Morse inequalities since the energy-functional $E:\Omega_{\{p\}\times\{q\}}(M)\longrightarrow\mathbb{R}$ is a Morse-function. ∎
Theorem 2 is Part (b) of Theorem 4.
Proof of Theorem 3.
As remarked in [7, p.117] the absence of orthogonal-tangent geodesic chords is an open condition with respect to the $C^{1}$-topology, hence $\mathcal{G}_{1}$ is an open subset. For a rotationally symmetric metric on the disc there are no orthogonal-tangent geodesic chords since any geodesic segment starting orthogonally from $\partial D^{n}$ will end orthogonally at the boundary $\partial D^{n}.$ Hence $\mathcal{G}_{1}$ is non-empty. Let $g_{0}$ be the Euclidean inner product on the $n$-disc $D^{n}\subset\mathbb{R}^{n},$ then let $g\in\mathcal{G}(D^{n})\longmapsto D(g)\in R^{+}$ be the smallest positive number such that $D(g)^{-1}g_{0}(v,v)\leq g(v,v)\leq D(g)g_{0}(v,v)$ for any tangent vector $v\in T_{x}D^{n},x\in D^{n}.$ Since the mapping $D:\mathcal{G}(D^{n})\longrightarrow\mathbb{R}^{+}$ is continuous the set $\mathcal{G}_{0}$ of Riemannian metrics for which all orthogonal geodesic chords of length $\leq\pi D(g)$ are non-degenerate and simple is an open and dense subset by Theorem 4(b), compare the argument in [19, Sec.6]. Then the intersection $\mathcal{G}_{2}=\mathcal{G}_{0}\cap\mathcal{G}_{1}$ contains all metrics $g$ which do not have an orthogonal-tangent geodesic chord and for which all orthogonal geodesic chords of length $\leq D(g)$ are simple and non-degenerate and do not intersect each other. Then we use the result by Giambò, Giannoni and Piccione [7, Thm.1.1] which ensures the existence of $n$ geometrically distinct orthogonal geodesic chords for a metric without orthogonal-tangent geodesic chords. ∎
If the boundary $\partial D^{n}$ is strictly convex there are no orthogonal-tangent geodesic chords, hence Theorem 3 implies Corollary 2.