### v2

# Geodesic loops and orthogonal geodesic chords without self-intersections

Abstract.
We show that for a generic Riemannian metric on a compact manifold of dimension $n\geq 3$ all geodesic loops based at a fixed point have no self-intersections. We also show that for an open and dense subset of the space of Riemannian metrics on an $n$-disc with $n\geq 3$ and with a striclty convex boundary there are $n$ geometrically distinct orthogonal geodesic chords without self-intersections. We use a perturbation result for intersecting geodesic segments of the author [23] and a genericity statement due to Bettiol and Giambò [4] and existence results for orthogonal geodesic chords by Giambò, Giannoni, and Piccione [9].
Key words and phrases:
orthogonal geodesic chords without self-intersections, perturbation of Riemannian metrics, non-degenerate orthogonal geodesic chord, orthogonal-tangent geodesic chord, generic Riemannian metrics
2020 Mathematics Subject Classification:
53C22, 58E10
1. Results
It is a well known result that on any compact smooth manifold $M$ with tangent bundle $TM$ endowed with a Riemannian metric $g$ there exists a non-trivial closed geodesic. If the manifold is not simply-connected a shortest non-contractible curve is a closed geodesic. In the simply-connected case this result is due to Lusternik and Fet [17], cf. [15, Sec.2.4]. We call a closed geodesic $c:S^{1}=\mathbb{R}/\mathbb{Z}\longrightarrow M$ prime if it is not the covering of a shorter closed curve, then the velocity field $c^{\prime}:S^{1}\longrightarrow TM$ defines an injective map. We call a closed geodesic simple resp. without self-intersections if the map $c:S^{1}\longrightarrow M$ is injective. In particular a simple closed geodesic is prime. The shortest noncontractible smooth curve on a manifold is a simple closed geodesic. On one hand any closed surfaces carries a simple closed geodesic, cf. [23, Rem. 5]. On the other hand it is an open question whether any Riemannian manifold $M$ of dimension $n=\dim(M)\geq 3$ carries a simple closed geodesic, cf. [2, Introduction]. The situation is quite different for a generic metric in dimensions $\geq 3.$ The author has shown recently that that for a generic Riemannian metric, resp. a generic reversible Finsler metric, on a compact manifold of dimension $n\geq 3$ all prime closed geodesics have no self-intersections, and they do not intersect each other, cf. [23, Thm.1].
One ingredient of the proof is the following perturbation result for intersecting geodesic segments: If $N\geq 2$ geometrically distinct geodesic segments $c_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,N$ parametrized by arc length intersect in the common midpoint $p=c_{j}(0)$ for sufficiently small $\eta>0$ then one can perturb away the intersection of the geodesic segments. For a precise statement see Lemma 2.
The other ingredient is the bumpy metric theorem for closed geodesics stating that for a generic Riemannian metric on a compact manifold all closed geodesics are non-degenerate in the sense of Morse theory. A closed geodesic $c:S^{1}\longrightarrow M$ is called non-degenerate if there is no non-trivial periodic Jacobi field along $c$ which is orthogonal to the velocity field $c^{\prime}.$ The statement is due to Abraham [1], proofs were given by Anosov [3], and White [27, Thm 2.2], see also the book by Moore [21, §2.7]. Hence the energy functional $E:\Lambda M\longrightarrow\mathbb{R},E(\sigma)=\frac{1}{2}\int_{0}^{1}g(\sigma^{\prime}(t),\sigma^{\prime}(t))\,dt$ on the Hilbert manifold $\Lambda M$ of closed curves $\sigma:S^{1}=[0,1]/\{0,1\}\longrightarrow M$ (called the free loop space) is a Morse-Bott function whose critical set is the disjoint union of orbits of closed geodesics under the canonical action of the orthogonal group $\mathbb{O}(2)$ in dimension $2$ acting on $S^{1}=\mathbb{R}/\mathbb{Z}=[0,1]/\{0,1\}.$ It also follows that for any $L>0$ there are only finitely many geometrically distinct closed geodesics with length $\leq L.$ It is a consequence of the perturbation result and the bumpy metric theorem that for a generic Riemannian metric on a compact manifold of dimension $n\geq 3$ all prime closed geodesics have no self-intersections, i.e. are simple, and they do not intersect each other, as shown by the author, cf. [23, Thm.1]. Note that on a surface with a non-degenerate closed geodesic with self-intersection it is not possible to perturb away the self-intersection.
Instead of considering closed resp. periodic geodesics one can study different boundary conditions for geodesic segments. In this paper we study first for points $p,q\in M$ geodesic segments $c:\mathcal{I}\longrightarrow M$ joining $p=c(0)$ and $q=c(1).$ Here we use as parameter space the unit interval $\mathcal{I}=[0,1]$ if $p\not=q$ and the circle $\mathcal{I}=S^{1}=\mathbb{R}/\mathbb{Z}=[0,1]/\{0,1\}$ in case of a geodesic loop, i.e. $p=q.$ Note that in the case of a geodesic loop $c:S^{1}\longrightarrow M$ only the map $c:[0,1]\longrightarrow M$ is smooth, since in general for a geodesic loop in contrast to a closed resp. periodic geodesic we have $c^{\prime}(0)\not=c^{\prime}(1).$ We call the geodesic segment $c:\mathcal{I}\longrightarrow M$ simple, resp. without self-intersection, if the map $c:\mathcal{I}\longrightarrow M$ is injective. We call two geodesic segments $c_{1},c_{2}:\mathcal{I}\longrightarrow M$ geometrically distinct, if $c_{1}(\mathcal{I})\not=c_{2}(\mathcal{I}).$ We say that two geometrically distinct geodesic segments $c_{1},c_{2}:\mathcal{I}\longrightarrow M$ do not have an interior intersection, if $c_{1}((0,1))\cap c_{2}((0,1))=\emptyset.$ It follows from a general result by Bettiol and Giambò presented in  [4, Theorem 5.10] that for a generic Riemannian metric on a compact manifold $M$ the geodesic segments, resp. geodesic loops, are non-degenerate in the sense of Morse theory, cf. Section 3. We show in this paper that we can use the perturbation result Lemma 2 and the genericity result of [4, Theorem 5.10] to show the following result:
Theorem 1.
Let $M$ be a compact manifold of dimension $n\geq 3.$ Let $p,q\in M$ be points in $M,$ where we also allow $p=q.$ For a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all geodesic segments joining $p$ and $q$ are non-degenerate and simple, and geometrically distinct geodesic segments do not have an interior intersection.
Note that Theorem 1 also shows that for points $p,q$ for a generic metric there is no closed geodesic passing through $p$ and $q,$ cf. Lemma 6. It even shows that for a generic metric any geodesic through $p$ and $q$ does not go through $p$ resp. $q$ again. In the particular case $p=q$ it follows from Theorem 1 that all geodesic loops for a generic metric are prime, i.e. there is no shorter geodesic loop, which is a part of the given geodesic loop, cf. [20]. Hence for a generic metric for any two geodesics $c,d:[0,1]\longrightarrow M,c(0)=d(0)=p,c(1)=d(1)=q$ we have either $c((0,1))=d((0,1))$ or $c([0,1])\cap d([0,1])=\{p,q\}.$
As a consequence of Theorem 1 we obtain the following
Corollary 1.
Let $M$ be a simply-connected and compact manifold of dimension $n\geq 3$ and $p,q\in M.$ For a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all geodesic segments joining $p$ and $q$ are non-degenerate and simple, and geometrically distinct ones do not have interior intersection. For the number $N_{g}(p,q;t)$ of geometrically distinct and simple geodesic segments joining $p$ and $q$ of length $\leq t$ we obtain $\liminf_{t\to\infty}(N_{g}(p,q;t)/t)>0.$
This is a quantitative version of Serre’s result [25, Ch.IV., Sec.7, Prop.14] giving a result concerning geometrically distinct geodesic segments between two fixed points. Therefore Corollary 1 is also satisfying from a geometric point of view, cf. the discussion in the doctoral thesis by Mentges [20, Sec.1.2] or by Tanaka [26, Problem B]. On a surface, i.e. for $n=2,$ we cannot expect to remove intersections of geodesic segments by a small perturbation. But for a point $p$ and a generic metric on a closed surface, all geodesic loops based at $p$ are prime, cf. Proposition 1.
Another boundary condition for geodesic segments is the following: Let $(M,g)$ be a compact Riemannian manifold with smooth boundary $\partial M$ and interior $\textrm{Int}(M)=M-\partial M$ then a geodesic chord $c:[0,1]\longrightarrow M$ is a geodesic segment with $c(0),c(1)\in\partial M$ and $c(t)\in\textrm{Int}(M)$ for all $t\in(0,1).$ It is called an orthogonal geodesic chord if $c$ meets the boundary $\partial M$ orthogonally. Orthogonal geodesic chords are also called free boundary geodesics, cf. for example [16] or [28]. A chord $c:[0,1]\longrightarrow M$ is called an orthogonal-tangent geodesic chord if it is tangent to the boundary at $c(0)$ and orthogonal to the boundary at $c(1).$ As a consequence of the genericity result by Bettiol and Giambò [4] we show:
Theorem 2.
Let $M$ be a compact manifold of dimension $n\geq 3$ with smooth boundary $\partial M.$ For a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all orthogonal geodesic chords are non-degenerate and have no self-intersections. In addition they all do not intersect each other.
See Section 3 for the definition of non-degeneracy for orthogonal geodesic chords.
Lusternik and Schnirelmann [18] have shown that a bounded domain in $\mathbb{R}^{n},n\geq 2$ with a smooth and convex boundary has at least $n$ double normals, i.e. lines intersecting the boundary orthogonally. This was generalized for a Riemannian metric, Bos [6, p.431] showed that on an $n$-disc $D^{n}$ with a Riemannian metric whose boundary $\partial D^{n}$ is strictly convex there are $n$ geometrically distinct orthogonal geodesic chords. Using a generalization of this result due to Giambò, Giannoni, and Piccione [9, Theorem 1.1] we obtain from Theorem 2:
Theorem 3.
For $n\geq 3$ let $\mathcal{G}(D^{n})$ be the space of Riemannian metrics on the $n$-disc $D^{n}$ with the strong $C^{k}$-topology with $k\geq 2.$ Let $\mathcal{G}_{1}=\mathcal{G}_{1}(D^{n})$ be the set of Riemannian metrics on $D^{n}$ for which there is no orthogonal-tangent geodesic chord. Then $\mathcal{G}_{1}$ is an open and non-empty subset of $\mathcal{G}(D^{n}).$
Furthermore there is an open subset $\mathcal{G}_{2}\subset\mathcal{G}_{1}$ whose closure $\overline{\mathcal{G}}_{2}$ coincides with the closure $\overline{\mathcal{G}}_{1},$ i.e. $\overline{\mathcal{G}}_{1}=\overline{\mathcal{G}}_{2},$ with the following property: For all Riemannian metrics $g\in\mathcal{G}_{2}$ there are $n$ geometrically distinct geodesic orthogonal chords, which are non-degenerate and have no self-intersection, and do not intersect each other.
Here we note that geometrically distinct geodesic chords cannot intersect at a boundary point. The boundary $\partial D^{n}$ is called strictly convex, if all principal curvatures $\kappa_{i},i=1,\ldots,n-1$ of the boundary with respect to the inward pointing unit normal, are positive, i.e. $\kappa_{i}>0,i=1,\ldots,n-1.$ In particular there is no geodesic in the disc touching the boundary $\partial D^{n}$ and hence there is no orthogonal-tangent geodesic chord. Therefore we obtain from Theorem 3 the following result in the spirit of Bos’ theorem [6, p.431]:
Corollary 2.
There is an open and dense subset $\mathcal{G}_{*}$ of the set $\mathcal{G}_{+}(D^{n})$ of Riemannian metrics on the $n$-dics $D^{n}$ of dimension $n\geq 3,$ for which the boundary $\partial D^{n}$ is strictly convex, with the following property: For all Riemannian metrics $g\in\mathcal{G}_{*},$ there are at least $n$ geometrically distinct orthogonal geodesic chords without self-intersections.
Furthermore, these geodesic chords do not intersect each other.
Using a curve shortening construction called disc flow on surfaces, Hass and Scott [13, Thm.4.2] have shown that a Riemannian $2$-disc with strongly convex boundary carries at least two geometrically distinct orthogonal geodesic chords without self-intersections. For the proof one can also use the curve shortening flow, as shown by Ko [16]. With a curve shortening flow, Li [19] presents a proof of the above cited result by Lusternik and Schnirelmann about the existence of $n$ double normals of a strictly convex compact domain in $\mathbb{R}^{n}.$
Acknowledgement. I am grateful to the referee for the helpful comments and suggestions.
2. Intersection of geodesic segments and perturbation
For the intersection points of geodesic segments we obtain the following result which is similar to the corresponding result for closed geodesics, cf. [23, Lem.2] resp. [4, Sec.2.3]. Recall from the Introduction that we denote a geodesic segment $c:\mathcal{I}\longrightarrow M$ joining $p=c(0),q=c(1)$ using the parameter space $\mathcal{I}=[0,1]$ if $p\not=q$ and $\mathcal{I}=S^{1}=[0,1]/\{0,1\}$ for $p=q.$
Lemma 1.
(a) Let $\gamma:\mathcal{I}\longrightarrow M$ be a geodesic segment, which is not the parametrization of a closed (periodic) geodesic of period $\leq 1.$ Then the set of interior double points, resp. interior self-intersection points
$DP_{1}(\gamma):=\{\gamma(t)\,;\,t\in(0,1),\#\gamma^{-1}(\gamma(t))\geq 2\}$
of the geodesic segment $\gamma$ is finite.
(b) Let $\gamma,\delta:[0,1]\longrightarrow M$ be geometrically distinct geodesic segments with $\gamma(0)=\delta(0)=p,\gamma(1)=\delta(1)=q.$ Then the set of interior intersection points
$I_{1}(\gamma,\delta):=\gamma((0,1))\cap\delta((0,1))$
is finite.
Remark 1.
(a) Note that in case (a) for a geodesic segment $\gamma:\mathcal{I}\longrightarrow M$ the initial point $c(0)$ is an interior self-intersection point, if there is a parameter $s\in(0,1)$ with $c(s)=c(0).$
(b) The geodesic segment $\gamma:\mathcal{I}\longrightarrow M$ has no interior self-intersection, resp. is called simple, if the map $\gamma$ is injective, i.e. if the set $DP_{1}(\gamma)$ of interior double points is empty. The geodesic segments $\gamma,\delta$ joining $p,q$ do not have interior intersection, if $I_{1}(\gamma,\delta)=\emptyset.$
Proof.
Let $\mathrm{inj}>0$ be the injectivity radius of $(M,g).$ We conclude from $\gamma(t)=\gamma(s)$ and $t<s$ that $(s-t)L(\gamma)\geq\mathrm{inj}.$ This implies that for all $t\in[0,1]:\#\gamma^{-1}(\gamma(t))\leq L(\gamma)/\mathrm{inj}+1.$
(a) If the set $DP_{1}(\gamma)$ is not finite, then there exist sequences $s_{j}<t_{j}\in[0,1]$ converging to $s^{*},t^{*}\in[0,1]$ with $s_{j}\not=s^{*},t_{j}\not=t^{*},(t_{j}-s_{j})L(\gamma)\geq\mathrm{inj}$ for all $j,$ $(t^{*}-s^{*})L(\gamma)\geq\mathrm{inj},$ and $\gamma(s_{j})=\gamma(t_{j})$ for all $j\geq 1.$ For $p_{1}=\gamma(t^{*})=\gamma(s^{*})$ set $v=c^{\prime}(t^{*}),w=c^{\prime}(s^{*}).$ The geodesic segment $\gamma$ is parametrized proportional to arc length, hence $\|v\|=\|w\|.$ Since $c(t_{j})=\exp_{p}((t_{j}-t^{*})v)=c(s_{j})=\exp_{p}((s_{j}-s^{*})w)$ we obtain for sufficiently large $j$ with $|t_{j}-t^{*}|\|v\|,|s_{j}-s^{*}|\|w\|<\mathrm{inj}$ that $(t_{j}-t^{*})v=(s_{j}-s^{*})w$ holds, i.e. $v=\pm w.$ If $v=w$ then $\gamma(t)=\gamma(t+(t^{*}-s^{*}))$ for all $t,$ i.e. $\gamma$ is a closed geodesic, which is not prime. If v=wv=-w we conclude $\gamma((t^{*}-s^{*})/2-t)=\gamma((t^{*}-s^{*})/2+t)$ for all $t,$ showing that $\gamma$ is not smooth at $(t^{*}-s^{*})/2.$
(b) The argument is similar to (a): If the set $I_{1}(\gamma,\delta)$ is infinite then there are sequences $s_{j},t_{j}\in[0,1]$ converging to $s^{*},t^{*}$ with $s_{j}\not=s^{*},t_{j}\not=t^{*}$ and $\gamma(s_{j})=\delta(t_{j})$ for all $j\geq 1.$ Let $p=\gamma(s^{*})=\delta(t^{*})$ and $v=\gamma^{\prime}(s^{*}),w=\delta^{\prime}(t^{*}).$ Since $\gamma$ and $\delta$ are geometrically distinct, we have $v/\|v\|\not=\pm w/\|w\|.$ Since $\gamma(t_{j})=\exp_{p}((t_{j}-t^{*})v)=\delta(s_{j})=\exp_{p}((s_{j}-s^{*})w)$ we conclude for sufficiently large $j$ with $|t_{j}-t^{*}|\|v\|,|s_{j}-s^{*}|\|w\|<\mathrm{inj}$ that $(t_{j}-t^{*})v=(s_{j}-s^{*})w$ holds, i.e. $v/\|v\|=\pm w/\|w\|,$ which is a contradiction. ∎
We obtain from the perturbation result [23, Lem.3]:
Lemma 2.
Let $(M,g)$ be a compact Riemannian manifold of dimension $n\geq 3$ and $\eta>0$ be a positive number, such that the injectivity radius $\mathrm{inj}$ satisfies $\mathrm{inj}\geq 4\eta.$ For a point $p\in M$ let $c_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,k$ be $k\geq 2$ geometrically distinct geodesic segments parametrized by arc length with common midpoint $p=c_{j}(0),j=1,\ldots,k.$
In any neighborhood $\mathcal{U}\ni g$ with respect to the strong $C^{k}$-topology with $k\geq 2,$ there is a Riemannian metric $\overline{g}\in\mathcal{U}$ coinciding with $g$ outside the geodesic ball $B_{p}(\eta)$ of radius $\eta$ centered at $p,$ carrying geodesic segments $\overline{c}_{j}:[-2\eta,2\eta]\longrightarrow M,j=1,\ldots,k$ parametrized by arc length, which coincide with $c_{j}$ outside $B_{p}(\eta).$
These geodesic segments are simple and do not intersect each other.
A similar proof as in Lemma [23, Lem.3] shows the following
Lemma 3.
Let $c:[0,1]\longrightarrow M$ be a geodesic loop with $p=c(0)=c(1)$ of the Riemannian metric $g$ on the differentiable manifold $M$ with $v=c^{\prime}(0)\in T_{p}M,w=c^{\prime}(1)\in T_{p}M.$ Then there exists a continuous one parameter perturbation $g_{s},s\in[0,s_{1}]\in\mathcal{G}(M)$ in the space of Riemannian metrics with the strong $C^{k}$-topology with $k\geq 2$ of the metric $g=g_{0},$ and smooth functions $s\in[0,s_{1}]\mapsto v(s),w(s)\in T_{p}M,$ such that $c_{s}:[0,1]\longrightarrow M$ is a geodesic loop with $p=c(0)$ of the Riemannian metric $g_{s}$ with $v(s)=c_{s}^{\prime}(0),w(s)=c_{s}^{\prime}(1),$ and $v(s),v$ (resp. $w(s),w$) are linearly independent for $s>0.$ The metric $g_{s}$ agrees with $g$ outside an arbitrarily small geodesic ball around $p.$
3. Geodesics with product boundary conditions
The following concept is introduced in [4, Sec.4], cf. also [12]. Let $M$ be a differentiable manifold. Let $\mathcal{P}\subset M\times M$ be a compact submanifold. This is called a general boundary condition on $M$ if for some Riemannian metric $g$ on $M$ the pseudo-Riemannian metric $\overline{g}=g\oplus(-g)$ on $M\times M$ induces a non-degenerate pseudo-Riemannian metric on $\mathcal{P},$ cf.  [4, Def.4.2]. It is shown in [4, Prop.4.1] that the subset $\mathcal{A}_{\mathcal{P}}\subset\mathcal{G}(M)$ of the set of Riemannian metrics on $M$ for which the restriction of the pseudo Riemannian metric $\overline{g}$ onto the submanifold $\mathcal{P}$ is non-degenerate, is an open subset. The second fundamental form of $\mathcal{P}$ in the normal direction $\eta\in T\mathcal{P}^{\perp}$ is given by
$S^{\mathcal{P}}_{\eta}(v,w)=\overline{g}\left(\nabla_{v}^{\overline{g}}\overline{w},\eta\right),$
cf. [4, p.346]. Here $\overline{\nabla}$ is the Levi-Civita connection on $M\times M$ with metric $\overline{g}.$ The vector field $\overline{w}$ is a smooth extension of $w$ to a smooth vector field along $\mathcal{P}.$ Since the restriction of $\overline{g}$ to $\mathcal{P}$ is non-degenerate we can identify the symmetric bilinear form with the self-adjoint shape operator $(S_{\eta}^{\mathcal{P}})_{p}:T_{p}\mathcal{P}\longrightarrow T_{p}\mathcal{P}$ with $\overline{g}((S_{\eta}^{\mathcal{P}})_{p}v,w)=S_{\eta}^{\mathcal{P}}(v,w)$ for all $v,w\in T_{p}\mathcal{P}.$ Let $H^{1}(I,M)$ be the set of all curves of Sobolev class $H^{1}$ in $M,$ i.e. of absolutely continuous curves with finite energy. Then the subspace
$\Omega_{\mathcal{P}}(M)=\left\{\gamma\in H^{1}(I;M)\,;\,(\gamma(0),\gamma(1))\in\mathcal{P}\right\}$
is a Hilbert manifold. The tangent space $T_{\gamma}\Omega_{\mathcal{P}}(M)$ can be identified with the completion of the space of smooth vector fields $v$ along $\gamma$ with $(\gamma(0),\gamma(1))\in T_{(\gamma(0),\gamma(1))}\mathcal{P}$ with respect to the induced metric $g_{1}=\langle.,.\rangle_{1}.$ It is defined on $T_{\gamma}\Omega_{\mathcal{P}}(M)$ by
$\langle v,w\rangle_{1}=\int_{0}^{1}g\left(v(t),w(t)\right)\,dt+\int_{0}^{1}g\left(\nabla v(t),\nabla w(t)\right)\,dt\,.$
A geodesic $\gamma:I\longrightarrow M$ is called a $(g,\mathcal{P})$-geodesic if $(\gamma^{\prime}(0),\gamma^{\prime}(1))\in T_{(\gamma(0),\gamma(1))}\mathcal{P}^{\perp}.$ Here the orthogonality is with respect to the metric $\overline{g}.$ It is shown in [4, Sec. 5] that a curve $\gamma$ is a critical point of the energy functional $E_{g}:\Omega_{\mathcal{P}}(M)\longrightarrow\mathbb{R},$
(1) $E_{g}(\gamma)=\frac{1}{2}\int_{0}^{1}g(\gamma^{\prime}(t),\gamma^{\prime}(t))\,dt$
if and only if it is a $(g,\mathcal{P})$-geodesic. The kernel of the index form is the space of Jacobi fields $Y\in T_{\gamma}\Omega_{\mathcal{P}}(M)$ such that
(2) $\left(\nabla Y(0),\nabla Y(1)\right)+S_{(\gamma^{\prime}(0),\gamma^{\prime}(1))}(Y(0),Y(1))\in T_{(\gamma(0),\gamma(1))}\mathcal{P}^{\perp}.$
[…]
A general boundary condition is called admissible if for every $g_{0}\in\mathcal{A}_{\mathcal{P}}$ there exists an open neighborhood $\mathcal{V}$ of $g_{0}$ in $\mathcal{A}_{\mathcal{P}}$ and a positive real number $a>0$ such that for all $g\in\mathcal{V}$ and all non-trivial $(g,\mathcal{P})$-geodesics $\gamma$ we have: $L_{g}(\gamma)\geq a,$ cf. [4, Def.4.5]. Here a non-trivial $(g,\mathcal{P})$ geodesic is a non-constant curve.
We restrict ourselves to the following case:
Definition 1.
Let $(M,g)$ be a complete Riemannian manifold of dimension $n$ and $N_{0},N_{1}\subset M$ connected and compact submanifolds (without boundary) of dimension $\dim N_{i}=n_{i}\in\{0,1,\ldots,n-1\},i=0,1.$ Then the product boundary condition is given as general boundary condition with $\mathcal{P}=N_{0}\times N_{1}.$
In particular a $(g,N_{0}\times N_{1})$-geodesic is a geodesic $c:[0,1]\longrightarrow M$ with $c(l)\in N_{l},c^{\prime}(l)\in T_{c(l)}^{\perp}N_{l},l=0,1.$ The product metric $g\oplus(-g)$ on $N_{0}\times N_{1}$ is non-degenerate, hence a product boundary condition is always non-degenerate, as defined in [4, Sec.4]. In case $\mathcal{P}=N\times N$ the $(g,N\times N)$-Jacobi fields are Jacobi fields $Y\in T_{\gamma}\Omega_{N\times N}$ with
(3) $\nabla Y(l)+S_{\gamma(l)}(Y(l))\in T_{\gamma(l)}N^{\perp}\,,l=0,1\,$,
they form the kernel of the index form. The $(g,N\times N)$-geodesic $\gamma$ is non-degenerate if the kernel is trivial, i.e. if there is no non-trivial $(g,N\times N)$-Jacobi field. Here $S_{x}$ denotes the shape operator of the submanifold $N\subset M$ with respect to a unit normal field $\nu,$ i.e.
$S_{x}(v)=\nabla_{v}\nu\,.$
And $\nabla$ is the Levi-Civita connection of $(M,g).$ If $N_{0}=\{p\},N_{1}=\{q\},$ then a $(g,\{p\}\times\{q\})$-Jacobi field along a $(g,\{p\}\times\{q\})$-geodesic segment $\gamma:[0,1]\longrightarrow M$ is a Jacobi field $Y$ vanishing at the end points, i.e. $Y(0)=0,Y(1)=0.$ Hence $q$ is not a conjugate point for $p$ along $\gamma$ if and only if there is no $(g,\{p\}\times\{q\})$-Jacobi field.
Lemma 4.
The following sets $\mathcal{P}$ are admissible product boundary conditions:
(a) Let $(M,g)$ be a compact Riemannian manifold (without boundary) of dimension $n\geq 2.$ If $p,q\in M$ are two points, where we also allow $p=q,$ then $\mathcal{P}=\{p\}\times\{q\}$ is an admissible product boundary condition.
(b) If $(M,g)$ is a compact Riemannian manifold of dimension $n\geq 2$ with smooth boundary $\partial M,$ then $\mathcal{P}=\partial M\times\partial M$ is an admissible product boundary condition.
(c) If $(M,g)$ is a compact Riemannian manifold (without boundary) of dimension $n\geq 2,$ and if $N\subset M$ is a compact submanifold of codimension one, then $\mathcal{P}=N\times N$ is an admissible product boundary condition.
Remark 2.
In case (b) we can consider the compact manifold $M$ with smooth boundary $\partial M$ endowed with a Riemannian metric $g$ as subset of a complete Riemannian manifold $(M_{1},g_{1}).$
Proof.
(a) We can choose $a=d(p,q)/2$ if $p\not=q,$ and $a=\mathrm{inj}/2$ for $p=q,$ here $\mathrm{inj}$ is the injectivity radius. Here we use that the distance as well as the injectivity radius are continuous functions on $\mathcal{G}^{r}(M)$ for $r\geq 2.$
(b) and (c). Here the normal exponential map of the submanifold $\partial M$ resp. $N$ is defined $\exp^{\perp}:\nu^{+}(\partial M)\longrightarrow M,$ resp. $\exp^{\perp}:\nu(N)\longrightarrow M.$ The set $\nu^{+}(\partial M)$ is the subset of the normal bundle of $\partial M$ of normal vectors pointing inward and $\nu(N)$ is the normal bundle of $N\subset M.$ Then there is a positive and continuous function $g\in\mathcal{G}^{r}(M)\mapsto\mathrm{inj}^{\perp}(g),$ which is also called the injectivity radius of the normal exponential map $\exp^{\perp},$ such that the restriction onto the normal vectors of length $\mathrm{inj}^{\perp}$ is a diffeomorphism onto its image. Any non-trivial $(g,\partial M\times\partial M)$-geodesic, resp. $(g,N\times N)$-geodesic $\gamma:[0,1]\longrightarrow M,$ satisfies $L_{g}(\gamma)\geq 2\mathrm{inj}^{\perp}(g).$
We will discuss the following examples in detail: If for some points $p,q\in M$ we have $N_{0}=\{p\},N_{1}=\{q\},$ then a $(g,\{p\}\times\{q\})$-geodesic is a geodesic segment joining $p$ and $q.$ Here we also allow the case $p=q,$ i.e. the case of geodesic loops based at $p.$ Another important case is $N_{0}=N_{1}=\partial N$ is a compact smooth boundary of an open subset $N$, then orthogonal geodesic chords are $(g,\partial{N}\times\partial{N})$-geodesics. And for a compact hypersurface $N\subset M$ i.e. a compact hypersurface $N$ we consider $(g,N\times N)$-geodesics, i.e. geodesics $c:[0,1]\longrightarrow M$ with $c(l)\in N,c^{\prime}(l)\in T_{c(l)}^{\perp}N,l=0,1.$
Lemma 5.
Let $M$ be a compact manifold of dimension $n\geq 2$ with smooth boundary $\partial M.$ Then there exists a compact manifold $M_{1}$ without boundary, such that $M$ is an open subset of $M_{1}$ and $\partial M$ is a smooth submanifold of $M_{1}.$
Proof.
It is a standard construction that one can add a cylindrical neck to the manifold $M$ with smooth boundary, i.e. there exists a Riemannian manifold $M_{0}$ with metric $g_{0}$ such that $M$ is an open submanifold and $g$ is the restriction of $g_{0}$ to $M.$ Furthermore, for some positive $\epsilon>0$ the normal exponential map of the hypersurface $\partial M$ defines an injective immersion $\phi:(t,x)\in[-\epsilon,0]\times\partial M\longmapsto M_{0}.$ The Riemannian metric on $M$ near $\partial M$ then has the form $dt^{2}+h_{t}$ for a smooth family $t\in[-\epsilon,0]\mapsto h_{t}\in\mathcal{G}(\partial M)$ of Riemannian metrics $h_{t}$ on $\partial M.$ Then we define the manifold $M_{0}$ as the disjoint union of $M$ and the cylinder $\epsilon,1]\times\partial M\ni(t,x)$ where we identify $(t,x)$ and $\phi(t,x)$ for all $(t,x)\in(-\epsilon,0]\times\partial M.$ We extend the Riemannian metric on $\epsilon,1]\times\partial M$ as $dt^{2}+h_{t}$ such that $h_{t}=h_{1}$ for all $t\in[2/3,1].$ Hence the set $\phi([2/3,1]\times\partial M)$ is a cylindrical end, which is isometric to the Riemannian product $2/3,1]\times(\partial M,h_{1}).$ This construction can be found for example in [10, Section 5].
We take two copies of $(M_{0},g_{0})$ and identify them along the boundary $\partial M_{0}.$ This way we obtain a compact Riemannian manifold $(M_{1},g_{1})$ which consists of two copies of $(M_{0},g_{0})$ without boundary, such that $\partial M_{0}$ is a totally geodesic submanifold, which is the fixed point set of an isometric involution interchanging the two copies of $(M_{0},g_{0}).$
Remark 3.
In the setting of Lemma 5 let $\phi:\mathcal{G}(M_{1})\longrightarrow\mathcal{G}(M)$ be the map defined by restricting the Riemannian metric $g$ on $M_{1}$ to the submanifold $M,$ i.e. $\phi(g)=g|M.$ The following statements are an immediate consequence of the definition of the strong topology, in particular of the definition of a strong basic neighborhood cf. [14, Section 2.1] applied to the space of Riemannian metrics on $M$ resp. $M_{1}.$ After fixing a Riemannian metric the space $\mathcal{G}(M)$ of Riemannian metrics of class $C^{k},k\geq 2$ on the manifold $M$ endowed with with the strong topology can be seen as an open subset in the Banach space of $C^{k}$-sections of the bundle of symmetric bilinear forms on the manifold $M.$ If $\mathcal{U}$ is an open and dense subset of $\mathcal{G}(M_{1})$ then $\phi(\mathcal{U})$ is an open and dense subset of $\mathcal{G}(M).$ And if $\mathcal{V}$ is a residual subset of $\mathcal{G}(M_{1}),$ then the restriction $\phi(\mathcal{V})$ is a residual subset of $\mathcal{G}(M).$
Lemma 6.
Let $L>0.$
(a) Let $(M,g)$ be a compact Riemannian manifold of dimension $n\geq 3$ and $p,q\in M.$ Then there is an open and dense set $\mathcal{G}^{*}[L]$ of Riemannian metrics such that for any $g\in\mathcal{G}^{*}[L]$ no geodesic segment of length $\leq L$ joining $p$ and $q$ is part of a closed geodesic.
(b) Let $(M,g)$ a compact Riemannian manifold of dimension $n\geq 2$ and let $N\subset M$ be a compact submanifold of codimension $1.$ Then there is an open and dense set $\mathcal{G}^{*}[L]$ of Riemannian metrics on $M$ such that for any $g\in\mathcal{G}^{*}[L]$ no $(g,N\times N)$-geodesic of length $\leq L$ is part of a closed geodesic.
Proof.
Assume that $\gamma:S^{1}=[0,1]/\{0,1\}\longrightarrow M$ is a closed geodesic. We show that the set $A(\gamma):=\{t\in S^{1};\gamma^{\prime}(t)\in T_{c(t)}^{\perp}N_{l}\},l=0,1$ is finite. Since the boundary condition is admissible there is a positive $a>0$ such that for $s,t\in A,s\not=t$ we have $|t-s|L(\gamma)\geq a.$ This implies that $A$ is a finite set.
For $L>0$ let $\mathcal{G}^{*}[L]$ be the set of Riemannian metrics for which all closed geodesics with energy $\leq L$ are non-degenerate and simple, cf. [23, Thm.1] such that no restriction onto a subinterval defines - up to linear reparametrization - a geodesic segment between $p$ and $q$ in case (a) resp. a $(g,N\times N)$-geodesic in case (b). This set is open, we have to show that it is a dense subset of $\mathcal{G}(M).$ A metric $g\in\mathcal{G}^{*}[L]$ has only finitely many geometrically distinct prime closed geodesics $c_{j}:S^{1}\longrightarrow M,j=1\ldots,k$ of energy $\leq L$ and there is an open subset $\mathcal{U}\subset\mathcal{G}^{*}[L]$ and continuous functions $g\in\mathcal{U}\mapsto v_{j}(g)\in TM$ such that the curve $t\in\mathbb{R}\mapsto c_{j}(t)=\Phi_{g}^{t}(v_{j}(g)),t\in S^{1}$ is a closed geodesic of $g$ and there are - up to geometric equivalence - no further closed geodesics of length $\leq L$ of the metric $g,$ cf. [3, p.12-13], resp. the Proof of Part (b) of Theorem 4. Here $\Phi^{t}_{g}:TM\longrightarrow TM,t\in\mathbb{R}$ denotes the geodesic flow, i.e. for $v\in T_{p}M,t\in\mathbb{R}\mapsto\Phi_{g}^{t}(v)\in TM$ is the periodic orbit in the tangent bundle whose projection on the manifold is the geodesic $c:\mathbb{R}\longrightarrow M$ with initial point $p=c(0)$ and initial direction $v=c^{\prime}(0).$
Hence the set $\widetilde{A}=\bigcup_{i=1}^{k}\{c_{j}(t)\,;\,t\in A(c_{j})\}$ is finite.
(a) We can perturb the metric in a sufficiently small neighborhood of the points $p$ and $q$ using Lemma 2 such that the perturbed closed geodesics of length $\leq L$ do not meet $p$ and $q.$
(b) We perturb the metric in a sufficiently small neighborhood of $p=c_{j}(t)\in\widetilde{A},t\in A(c_{j})$ such that $c_{j}$ remains a closed geodesic but $c_{j}^{\prime}(t)\not\in T_{c_{j}(t)}^{\perp}N_{l}$ with respect to the perturbed metric.
This is possible as the following Lemma shows. ∎
Lemma 7.
Let $D^{n-1}_{2\eta}=\{x\in\mathbb{R}^{n-1}\,;\,\|x\|\leq 2\eta\}$ be the Euclidean (n1)(n-1)-disc of radius $2\eta$ for $\eta>0.$ On $(t,x)=(t,x_{1},\ldots,x_{n-1})\in[-2\eta,2\eta]\times D^{n-1}_{2\eta}$ a Riemannian metric is given by
$dt^{2}+\sum_{i,j=1}^{n-1}g_{ij}(t,x)dx_{i}dx_{j},$
hence $\gamma(t)=(t,0,\ldots,0),t\in[-2\eta,2\eta]$ is a geodesic of $g$ parametrized by arc length orthogonal at $p=\gamma(0)=(0,\ldots,0)$ to the hypersurface $\Sigma:=\{0\}\times D^{n-1}_{2\eta}.$ Then in any neighborhood of the metric $g$ there is a metric $\overline{g}$ for which $\gamma$ is still a geodesic parametrized by arc length and which is not orthogonal to $\Sigma$ at $p$ with respect to $\overline{g}.$
Proof.
Choose a smooth bump function $f:[-2\eta,2\eta]\longrightarrow[0,1]$ with $f(t)=0$ for $|t|\geq\eta,$ $f(t)=1$ for $|t|\leq\eta/2.$ Let $s=t+\epsilon f(t)x_{1}.$ For sufficiently small $\epsilon>0$ with inverse function $t_{\epsilon}(s,x_{1})$ in the coordinates $s,x_{1},\ldots,x_{n-1}$ the metric is of the form
$\frac{ds^{2}+2\epsilon f(t)dsdx_{1}+\epsilon^{2}f^{2}(t)dx_{1}^{2}}{(1-\epsilon f^{\prime}(t))^{2}}+\sum_{i,j=1}^{n-1}g_{ij}(t_{\epsilon}(s,x_{1}),x_{1},\ldots,x_{n-1})dx_{i}dx_{j}\,.$
Then for sufficiently small $\epsilon>0$ the metric $\overline{g}$ given by
$ds^{2}+\sum_{i,j=1}^{n-1}g_{ij}(t_{\epsilon}(s,x_{1}),x_{1},\ldots,x_{n-1})dx_{i}dx_{j}$
is sufficiently close go $g$ in the $C^{k}$-topology for $k\geq 2$ and the the curve $x_{1}=\ldots=x_{n-1}=0$ coincides with $\gamma$ and is a geodesic of $\overline{g}$ parametrized by arc length. At the point $p=\gamma(0)$ the curve is orthogonal to the hypersurface $s=0$ but not to the hypersurface $\Sigma=\{t=0\}.$
With the help of Lemma 6 we can prove the following main result, from which we can derive the theorems stated in the Introduction:
Theorem 4.
Let $(M,g)$ be a compact Riemannian manifold of dimension $n\geq 3.$
(a) Let $M$ be a compact differentiable manifold of dimension $n\geq 3.$ Let $p,q\in M.$ Then for a $C^{k}$-generic Riemannian metric with $k\geq 2$ on $M$ all geodesic segments between $p$ and $q$ are non-degenerate and simple, and any geometrically distinct geodesic segments between $p$ and $q$ do not have interior intersection.
(b) Let $M$ be a compact differentiable manifold of dimension $n\geq 3$ with a smooth boundary $\partial M.$ Then for a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all orthogonal geodesic chords are non-degenerate and simple, and any two geometrically distinct orthogonal geodesic chords do not intersect.
(c) Let $M$ be a compact differentiable manifold (without boundary) of dimension $n\geq 3$ and let $N\subset M$ be a compact submanifold of codimension $1.$ Then for a $C^{k}$-generic metric on $M$ with $k\geq 2$ all $(g,N\times N)$-geodesics are non-degenerate and simple, and any two geometrically distinct $(g,N\times N)$-geodesics do not intersect.
Proof.
We first consider the three cases (a), (b), (c) at the same time, i.e. we set $\mathcal{P}=N_{0}\times N_{1}=\{p\}\times\{q\},$ or $\mathcal{P}=\partial{M}\times\partial{M},$ or $\mathcal{P}=N\times N.$
It follows from Lemma 6 and from the result [4, Thm.5.1], cf. also [4, p.347], that for $L>0$ the set $\mathcal{G}_{\mathcal{P}}(M)[L]$ of metrics $g\in\mathcal{G}_{\mathcal{P}}(M)$ for which all $(g,\mathcal{P})$-geodesics of length $\leq L$ are nondegenerate and are not part of a closed geodesic is an open and dense subset of $\mathcal{G}_{\mathcal{P}}=\mathcal{G}_{\mathcal{P}}(M).$ This holds since in all three cases (a),(b),(c) the boundary conditions $\mathcal{P}$ are non-degenerate and admissible, cf. Lemma 4.
Here one has to note that in case (b) one has to consider the differentiable manifold $M$ to be a subset of a compact manifold $M_{1}$ without boundary, cf. Lemma 5. Then one can apply [4, Thm.5.1] and conclude that the set $\mathcal{G}_{\partial M\times\partial M}(M_{1})[L]$ of metrics $g\in\mathcal{G}(M_{1})$ for which all $(g,\partial M\times\partial M)$-geodesics of length $\leq L$ are non-degenerate is an open and dense subset of $\mathcal{G}_{\partial M\times\partial M}(M_{1}).$ Then the set $\mathcal{G}_{\partial M\times\partial M}(M)$ contains all restrictions $g|M$ of metrics $g\in\mathcal{G}_{\partial M\times\partial M}(M_{2})[L],$ and is therefore also dense in $\mathcal{G}(M),$ cf. Remark 3.
For any $L>0$ we define the set $\mathcal{G}_{\mathcal{P}}^{*}(M)[L]$ of metrics $g\in\mathcal{G}^{L}(M)$ for which all $(g,\mathcal{P})$-geodesics of length $\leq L$ are nondegenerate and simple and such that for any pair $\gamma,\delta$ of geometrically distinct $(g,\mathcal{P})$-geodesics of length $\leq L$ in case of $\mathcal{P}=\{p\}\times\{q\}$ the interior intersection set $I_{1}(\gamma,\delta)$ is empty, and in the other two cases the intersection $\gamma(I)\cap\delta(I)=\emptyset.$
We conclude from [12, Thm.2.4] that the energy functional $E_{g}:c\in\Omega_{\mathcal{P}}(M)\longmapsto E_{g}(c)\in\mathbb{R}$ satisfies the Palais-Smale condition. We can apply this result since $M$ is compact, hence any Riemannian metric on $M$ is complete.
Hence for $g\in\mathcal{G}_{\mathcal{P}}(M)[L]$ the restriction $E_{g}:\Omega_{\mathcal{P}}(M)\cap\{E\leq L\}\longrightarrow\mathbb{R}$ is a Morse function with only finitely many critical points, i.e. there are only finitely many geometrically distinct $(g,\mathcal{P})$-geodesics $c_{j},j=1,\ldots,k$ of energy $\leq L.$
The self-adjoint operator $H_{c}$ representing the hessian $d^{2}E_{g}(c)$ at a $(g,\mathcal{P})$-geodesic $c$ is of the form $H_{c}=\mathbbm{1}+k_{c},$ i.e. the sum of the identity and a compact operator $k_{c}.$ A proof can be found in[15, Lem.2.5.2] or in [4, Sec.5.1], where one can set $g_{0}=g_{R},$ hence $A=\mathbbm{1}.$ Since 0 is not an eigenvalue the inverse $H_{c}^{-1}$ is a bounded linear operator. Therefore we can use the implicit function theorem for Banach manifolds, cf. [7, Thm.15.1] and conclude: There is a neigbhourhood $\mathcal{U}$ of $g$ and continuous functions $g\in\mathcal{U}\longmapsto c_{j}[g]\in\Omega_{\mathcal{P}}(M),j=1,\ldots,k$ such that $c_{j}[g]$ is a $(g,\mathcal{P})$-geodesic of $g\in\mathcal{U}.$
One can choose the open neighborhhod $\mathcal{U}$ sufficiently small such that there are no further $(g,\mathcal{P})$-geodesics of $g$ with $E_{g}(\gamma)\leq L,$ cf. the argument in [3, p.12-13].
Since the geodesics $c_{j}$ are not part of a closed geodesic we conclude from Lemma 1: The union $DIP_{1}(L)$ of the double points $DP_{1}(c_{j}),j=1,\ldots,k$ and the intersection points $I(c_{j},c_{l}),j\not=l,j,k\in\{1,\ldots,k\}$ is finite.
Let $x\in DIP_{1}(L).$ In cases (b)(c) it follows that if $x=c_{j}(t)$ for some $j=1\ldots,k$ then $t\in(0,1).$ In case (b) a geodesic chord $c_{j}:[0,1]\longrightarrow M$ satisfies $c_{j}(t)\not\in\partial M$ for all $t\in(0,1),$ hence $DIP_{1}(L)\cap\partial M=\emptyset.$
In case (c) if $x\in DIP_{1}(L)$ satisfies $x=c_{j}(0)=c_{j}(t)$ for some $t\in(0,1]$ then $c_{j}$ lies on a closed geodesic which is not possible by assumption. If $x=c_{j}(0)=c_{k}(t)$ for $j\not=k$ then $t\not\in\{0,1\}$ since otherwise the geodesics $c_{j},c_{k}$ lie on the same geodesic line $\gamma:\mathbb{R}\longrightarrow M$ since $N$ has codimension one. If $x=c_{j}(0)=c_{i}(t)$ for $i,j=1,\ldots,k,0<t<1$ then we can use the perturbation result Lemma 2 and obtain a Riemannian metric for which the geodesic segment $c_{j}$ in the neighborhood remains unchanged and the geodesic segment $c_{k}$ in the neighboorhood is perturbed such that $c_{j}$ and $c_{k}$ do not intersect in the neighborhood of $x.$
Now we consider case (a).
We first assume $p=q.$ For a Riemannian metric $g\in\mathcal{G}_{\{p\}\times\{p\}}(M)[L]$ for some $L>0$ there are only finitely many geometrically distinct and non-degenerate geodesic loops $\gamma_{1},\ldots,\gamma_{N}$ of length $\leq L.$ Assume that - after a possible renumbering - the geodesic loops $\gamma_{1},\ldots,\gamma_{m},m\leq N$ are the prime geodesic loops of length $\leq L.$ Here a geodesic loop $\gamma:[0,1]/\{0,1\}\longrightarrow M$ is prime, if there is no $t_{1}\in(0,1)$ with $p=c(t_{1}).$ Let $v_{j,l}=\gamma_{j}^{\prime}(l),j=1,\ldots,m,l\in\{0,1\}$ be the sequence of tangent vectors of the prime geodesic loops at $p.$ We use the perturbation result Lemma 3 which shows that there is an open and dense subset $\mathcal{G}_{p}[L]\subset\mathcal{G}_{\{p\}\times\{p\}}(M)[L]$ for which all geodesic loops at $p$ of length $\leq L$ are prime. Note here that if $m<N,$ then the geodesic loops $\gamma_{m+1},\ldots,\gamma_{N}$ are not prime, but their prime parts are loops $\gamma_{j},1\leq j\leq m.$ Since after the perturbation these prime geodesic loops cannot be concatenated to produce a non-prime geodesic loop the perturbed loops $\gamma_{j}[g],j\geq m+1$ have to be prime, too.
In case $p\not=q$ we consider the space $\mathcal{G}_{p,q}[L]=\mathcal{G}_{p}[L]\cap\mathcal{G}_{q}[L]\cap\mathcal{G}_{\{p\}\times\{q\}}(M)[L].$ Hence for $g\in\mathcal{G}_{p,q}[L]$ all geodesic loops at $p$ and $q$ of length $\leq L$ are non-degenerate and prime and there are only finitely many geometrically distinct geodesic loops at $p$ and $q.$ And all geodesic segments joining $p$ and $q$ of length $\leq L$ are non-degenerate and there are only finitely many geometrically distinct ones. Hence the set of tangent vectors of geodesic loops at $p$ resp. $q$ of length $\leq L$ as well as the set of tangent vectors of prime geodesic segments joining $p$ and $q$ of length $\leq L$ at $p$ and $q$ is finite. Therefore we can use the perturbation lemma 3 a finite number of times to obtain a metric for which all geometric segments joining $p$ and $q$ are prime. Here the observation is that a geodesic segment between $p$ and $q$ which is not prime contains either distinct prime geodesic segments joining $p$ and $q$ or contains prime geodesic segments and geodesic loops at $p$ or at $q.$
Therefore this argument shows that for points $p,q$ (here we also allow $p=q$) the set $\mathcal{G}^{\prime}_{\{p\}\times\{q\}}[L]$ of Riemannian metrics for which all geodesic segments of length $\leq L$ joining $p$ and $q$ are non-degenerate and prime is an open and dense subset. If now $x\in DIP_{1}(L),x\not\in\{p,q\}$ we can choose $\eta\in(0,\mathrm{inj}/3)$ such that the following holds: If a geodesic $c_{j}$ for some $j=1,\ldots,k$ enters the geodesic ball $B_{p}(2\eta)$ around $p\in DIP(L)$ with radius $2\eta,$ i.e. $c_{j}(s)\in B_{p}(2\eta),$ then there is $s_{1}$ with $c_{j}(s_{1})=p$ and $d(c_{j}(s),c_{j}(s_{1}))=|s-s_{1}|\|c_{j}^{\prime}\|\leq 2\eta.$ Hence one can use the perturbation result [23, Lem.3] resp. Lemma 2 and perturb the metric in the ball $D_{p}(2\eta)$ for all finitely many points in $DIP_{1}(L)$ such that we obtain a metric $g_{1}\in\mathcal{G}_{\{p\}\times\{q\}}^{*}(M)[L]$ in any arbitrary small neighborhood of the original metric. ∎
4. Proofs of the statements in section 1
Theorem 1 is Part (a) of Theorem 4.
Proof of Corollary 1.
Since $M$ is simply-connected and compact the following holds: There is a number $l\in\{2,\ldots,n-1\},$ such that the sequence $b_{2lk}=\dim H_{2lk}\left(\Omega_{\{p\}\times\{q\}}(M);\mathbb{Q}\right),k\geq 1$ satisfies $b_{2lk}\geq 1.$ It was shown by Gromov [11] and Paternain [22, Thm.5.1] that there is a constant $C(g)>0$ such that the critical value $\mathrm{cr}(w_{k})$ of a non-trivial homology class $w_{k}\in H_{2lk}\left(\Omega_{\{p\}\times\{q\}}(M);\mathbb{Q}\right)$ satisfies $\mathrm{cr}(w_{k})\leq C(g)k$ for all $k\geq 1.$ By Theorem 1 we obtain for a generic metric $g$ a sequence $c_{k}$ of simple and non-degenerate geodesic segments joining $p$ and $q$ of index $2lk$ and length $\leq C(g)k.$ This follows from the Morse inequalities since the energy-functional $E:\Omega_{\{p\}\times\{q\}}(M)\longrightarrow\mathbb{R}$ is a Morse-function. ∎
Note that a geodesic loop $c:[0,1]/\{0,1\}\longrightarrow M$ with $p=c(0)$ is prime, if there is no $t\in(0,1)$ with $p=c(t).$ Going through the Proof of Theorem 4 and of Corollary 1 we find that in case of a surface, i.e. for $n=2,$ we obtain the following result. Here we essentially use the Perturbation Lemma 3:
Proposition 1.
Let $M$ be a compact surface, i.e. $n=\dim M=2$ and let $p\in M$ be a point.
  • (a)
    For a $C^{k}$-generic Riemannian metric on $M$ with $k\geq 2$ all geodesic loops based at $p$ are non-degenerate and prime.
  • (b)
    For a $C^{k}$-generic Riemannian metric on the $2$-sphere $S^{2}$ the number $N_{g}(p;t)$ of geodesic loops based at $p$ satisfies $\liminf_{t\to\infty}(N_{g}(p;t)/t)>0.$
Theorem 2 is Part (b) of Theorem 4.
Proof of Theorem 3.
As remarked in [9, p.117] the absence of orthogonal-tangent geodesic chords is an open condition with respect to the $C^{1}$-topology, hence $\mathcal{G}_{1}$ is an open subset. For a rotationally symmetric metric on the disc there are no orthogonal-tangent geodesic chords since any geodesic segment starting orthogonally from $\partial D^{n}$ will end orthogonally at the boundary $\partial D^{n}.$ Hence $\mathcal{G}_{1}$ is non-empty. Let $g_{0}$ be the Euclidean inner product on the $n$-disc $D^{n}\subset\mathbb{R}^{n},$ then let $g\in\mathcal{G}(D^{n})\longmapsto D(g)\in R^{+}$ be the smallest positive number such that $D(g)^{-1}g_{0}(v,v)\leq g(v,v)\leq D(g)g_{0}(v,v)$ for any tangent vector $v\in T_{x}D^{n},x\in D^{n}.$ Since the mapping $D:\mathcal{G}(D^{n})\longrightarrow\mathbb{R}^{+}$ is continuous the set $\mathcal{G}_{0}$ of Riemannian metrics for which all orthogonal geodesic chords of length $\leq\pi D(g)$ are non-degenerate and simple is an open and dense subset by Theorem 4(b), compare the argument in [24, Sec.6]. Then the intersection $\mathcal{G}_{2}=\mathcal{G}_{0}\cap\mathcal{G}_{1}$ contains all metrics $g$ which do not have an orthogonal-tangent geodesic chord and for which all orthogonal geodesic chords of length $\leq D(g)$ are simple and non-degenerate and do not intersect each other. Then we use the result by Giambò, Giannoni and Piccione [9, Thm.1.1] which ensures the existence of $n$ geometrically distinct orthogonal geodesic chords for a metric without orthogonal-tangent geodesic chords. ∎
If the boundary $\partial D^{n}$ is strictly convex there are no orthogonal-tangent geodesic chords, hence Theorem 3 implies Corollary 2.